### v1

# Self-Similar Structure of Loop Amplitudes and Renormalization

Abstract
We study the self-similar structure of loop amplitudes in quantum field theory and apply it to amplitude generation and renormalization. A renormalized amplitude can be regarded as an effective coupling that recursively appears within another loop. It is best described as a coupling in the effective action. It is a scale-dependent, finite, parametrically small and observable quantity appearing in the $S$-matrix. Replacing a constant coupling with a loop amplitude provides a systematic method of generating high-order loop amplitudes, guaranteeing no subamplitude divergence. This approach also provides an alternative bottom-up method to the traditional top-down recursive renormalization of general amplitudes.
1 Introduction
Renormalization is about understanding the self-similarity of a system across varying scales. Its classic illustration is the block spin transformation, as discussed in Kadanoff’s seminal work [1]. In that, scaling up to include more spins allows for averaging them into a single spin, leading to a modified theory with the same formal structure but varying scale-dependent parameters. The scale is specified by temperature or site size. These modified parameters, known as “renormalized” physical parameters, are pivotal in renormalization, capturing its essence with varying scales.
In this work, we employ the same idea to describe the processes involving elementary fields in quantum field theory (QFT). Physical parameters acquire scale dependence via renormalization and have a recursive structure. A renormalized loop amplitude exhibits a recursive structure wherein it functions as a coupling constant within other amplitudes or recursively within itself. This recursive property enables the sum of such amplitudes to behave as an effective coupling analogous to the coupling observed in a free field theory.
Although the idea of self-similarity is well-established in statistical physics, it has yet to be fully utilized in particle physics. An exception is the famous running of the gauge couplings, which experiments have verified and led to the understanding of asymptotic freedom. The difficulty has been that in QFT, renormalization is mixed with the removal of infinities. Although loop corrections in QFT give rise to ultraviolet (UV) divergences, treating divergence is not the essence of renormalization.
The loop expansion is of a quantum nature because it accompanies powers of its characterizing Planck constant $\hbar$ [2]. Moreover, the recursive structure of the loop amplitudes from renormalizable couplings, in quantum field theory, is the key to understanding self-similarity. The self-similarity of theory is reflected in perturbative corrections to physical parameters. Different scattering processes involving the same asymptotic states are indistinguishable and thus are superposed. As a result, amplitudes at various loop orders are added up or exchangeable. Also, inside an amplitude, another loop amplitude can be nested, and vice versa.
Recent developments have rediscovered that a renormalized quantity expressed in terms observables is a meaningful physical one [3, 4, 5, 6]. Crucially, in experimental settings, these renormalized parameters are what we observe, as they can be measured only by interactions with other particles and with our measurement apparatus, with finite external momenta [3, 4]. This highlights a fundamental aspect: we should neither deal with the bare parameters in the defining Lagrangian nor the loop correction only, which are not separately observable in experiments. Suppose we use the physical parameters at a scale and the relative loop correction amplitude. In that case, we can define the same theory using the same number of finite parameters [5, 6], and the theory is finite and perturbatively well-defined.
This recursive structure is well-encoded in the effective parameters, such as loop-corrected mass and $n$-point couplings, which are the coefficients of the effective action. They span Green’s function and are the ones that are directly compared with the experimental data. The use of effective coupling eases the task of generating fully-renormalized higher-order amplitude systematically. The recursive generation of higher order amplitude is developed [7, 8, 9, 10, 11, 12]. Polchinski’s exact renormalization group [13] provides another method to reduce higher-order loops by integrating out high-loop operators by shrinking the loops. It showed that the vector space of the operator coefficients, in the $\phi^{4}$ theory, flows into three dimensional one for kinetic, mass, and dimension four terms.
In this work, we generate the higher-loop amplitudes and renormalize them simultaneously. If we use the renormalized physical coupling to calculate the amplitude, we do not have subdiagram divergence; thus, the renormalization is automatic. It is crucial to deal with the renormalized propagator when doing this.
Another important application is the full renormalization including the subamplitudes. It was completed by the program by Bogoliubov, Parasiuk, Hepp and Zimmermann (BPHZ) [14, 15, 16], extended for the massless fields [18, 19] (see also[20, 22, 21, 23, 24] for pedagogical introduction). This is a top-down approach that, from a given amplitude, tracks possible divergence of subamplitudes recursively. The Bogoliubov relation gives the relation between an amplitude and its subamplitude, and Zimmermann’s forest formula provides the explicit solution. Our method provides an alternative bottom-up method and helps us understand how the forest formula works.
This paper is organized as follows. First, in Section 2, after taking examples in the famous $\phi^{4}$-theory using renormalization “without counterterms.” [6]. The regularization independence of the effective parameters enables us to use loop-corrected coupling in place of tree-level coupling. We take an example of generating two-loop correction of the propagator and the quartic coupling, using “one-loop correction” involving “one-loop corrected effective coupling.” This replacement is well-described in terms of coupling in the effective action. In Section 3, we estabilish the concepts appropriate to describe the self-similar nature of the amplitude and develop tools for investigating it. In Section 4, we make use of this to generate general loop diagrams at arbitrary orders by iteration. Conversely, we can prove the renormalizability of a given diagram of arbitrary order, which is done in Section 5.
Notation
We usually denote a single $n$-point amplitude $A^{(n)}$ and the complete sum of $n$-point effective coupling as $\Gamma^{(n)}$. The subscript of a physical quantity, for instance, $l$ in $\tilde{\Sigma}_{l}(p^{2})$, denotes the loop-correction order, not the power of the coupling for the perturbation. We make a convention to use the round bracket for accumulated loop orders, while we do not use the bracket for individual orders. That is, $\Gamma^{(n)}_{(L)}({\underline{p}})=\sum_{l=0}^{L}\Gamma^{(n)}_{l}({\underline{p}}).$ If there is no subscript and does not depend on the external momenta ${\underline{p}}$, it is a zeroth order value, e.g., $\Gamma^{(n)}=\Gamma^{(n)}_{0}$. We usually choose such values as the “physical” ones observable in the scattering experiments. For the renormalization of $\Gamma^{(n)}_{l}({\underline{p}})$, we use a bar as in $\overline{\Gamma}^{(n)}_{l}({\underline{p}})$.
2 Two-loop amplitudes as effective one-loop ones
We present our main idea in the simplest case: we show two-loop amplitudes in $\phi^{4}$ theory are formally reduced to one-loop ones with “effective coupling.”
2.1 New Old Renormalization
We first identify a physical parameter that can be compared with experimental data. In this paper, we are mainly interested in the $\phi^{4}$-theory, which appears as the Higgs sector in the Standard Model. The Lagrangian density is
${\cal L}=\frac{1}{2}\partial_{\mu}\phi\partial^{\mu}\phi-\frac{1}{2}m_{B}^{2}\phi^{2}-\frac{6\lambda_{B}}{24}\phi^{4}.$ (1)
Here, the mass $m_{B}^{2}$ and the quartic coupling $\lambda_{B}$ are bare parameters defining the theory. They are defined in the absence of further interactions so that they are never directly observable. They receive quantum corrections so that the observed parameters are modified. Such corrections are divergent and we extract finite quantities by renormalization. We briefly discuss renormalization “without counterterms,” whose details are found in [6].
Consider the quartic coupling $\lambda_{B}$. We have a one-loop corrected coupling
$\lambda_{(1)}(p_{1},p_{2},p_{3},p_{4})=\lambda_{B}Z^{2}+6\lambda^{2}\left[V(s)+V(t)+V(u)\right],$ (2)
where
$iV(p^{2})=\frac{1}{2}\int\frac{d^{4}k}{(2\pi)^{4}}\frac{i}{(p+k)^{2}-m^{2}+i\epsilon}\frac{i}{k^{2}-m^{2}+i\epsilon},$ (3)
In what follows, we will omit the Feymann prescription $i\epsilon$. The factor $Z^{2}$ is reserved for the field renormalization that we see shortly. We have three channels described by the Mandelstam variables
$s=(p_{1}+p_{2})^{2},\quad t=(p_{1}+p_{3})^{2},\quad u=(p_{1}+p_{4})^{2},\quad s+t+u=4m^{2}.$ (4)
Figure 1: One-loop correction for the quartic coupling in $\phi^{4}$ theory. The first line after the renormalization is the effective coupling $\lambda_{1}({\underline{p}})$. We denote the corresponding weighted Feynman diagram as the same vertex and put the loop order below it as the one on each left-hand side.
We should be able to make these results to contact observations. We impose the renormalization condition at a reference scale $(s,t,u)=(s_{0},t_{0},u_{0})$ such that the scattering matrix element ${\cal M}$ for four scalars becomes the coupling $\lambda$
$-6\lambda_{(1)}\left(s_{0},t_{0},u_{0}\right)={\cal M}\left(\text{4 scalars}\right)\equiv-6\lambda+{\cal O}(\hbar^{2}\lambda^{3}).$ (5)
This gives us the relation between $\lambda$ and $\lambda_{B}Z^{2}$
$6\lambda=6\lambda_{(1)}\left(s_{0},t_{0},u_{0}\right)=\lambda_{B}Z^{2}+36\lambda^{2}\left[V(s_{0})+V(t_{0})+V(u_{0})\right]+{\cal O}(\hbar^{2}\lambda^{3}).$ (6)
These two equations enable us to remove the bare parameter $\lambda_{B}Z^{2}$ to give
$\begin{split}6\lambda_{(1)}(p_{1},p_{2},p_{3},p_{4})&=6\lambda+36\lambda^{2}\left[V(s)+V(t)+V(u)-V(s_{0})-V(t_{0})-V(u_{0})\right]\\ &=6\lambda+\frac{36\lambda^{2}}{32\pi^{2}}\int_{0}^{1}dx\log\frac{m^{2}-x(1-x)s}{m^{2}-x(1-x)s_{0}}+(s\leftrightarrow t)+(s\leftrightarrow u).\end{split}$ (7)
That is, ours is the relative coupling in the unit of the reference coupling $\lambda$ defined in (5).
The renormalized coupling (7) is regularization-independent because it is finite. We may verify that any choice of regularization gives the same result. Note that each $V(s)$ is divergent, but the combination $V(s)-V(s_{0})$ is not, and hence $6\lambda_{(1)}(p_{1},p_{2},p_{3},p_{4})$ is finite as well. The latter is parametrically small and admits perturbative expansion.
It is a well-defined and observable physical coupling. We can define the equivalent theory using this physical coupling $\lambda$ instead of the bare coupling $\lambda_{B}$ thanks to the relation (6). This is possible because we are dealing with the effective field theory, defined below a certain scale and the parameter in the Lagrangian is just a parametrization of our ignorance. It is to be matched with the counterpart in the UV completion or simply the observables.
Next, the mass is similarly corrected by the self-energy $\tilde{\Sigma}(p^{2})$. The quadratic interaction becomes
$p^{2}-m_{B}^{2}-\tilde{\Sigma}(p^{2})$ (8)
We need a reference mass for expanding the theory. One natural choice is the pole mass $m$, defined as
$\left[p^{2}-m_{B}^{2}-\tilde{\Sigma}(p^{2})\right]_{p^{2}=m^{2}}=0,$ (9)
Using this, we can eliminate the bare mass and express the corrected quadratic interaction (8) as $p^{2}-m^{2}-\tilde{\Sigma}(p^{2})+\tilde{\Sigma}(m^{2})$. We define the propagator as its inverse
$D^{\prime}(p^{2})=\frac{i}{p^{2}-m^{2}-\tilde{\Sigma}(p^{2})+\tilde{\Sigma}(m^{2})},$ (10)
whose pole is located indeed at $p^{2}=m^{2}$, justifying the name of the reference mass. Taking into account the correction, the residue of the propagator at this point is different from that of the free field
$p^{2}\to m^{2}:D^{\prime}(p^{2})\to\frac{iZ}{p^{2}-m^{2}}+{\cal O}((p^{2}-m^{2})^{2}),\quad Z^{-1}=1-\frac{d\tilde{\Sigma}}{dp^{2}}(m^{2}),$ (11)
which comes from the next leading order expansion from $\tilde{\Sigma}(p^{2})$.
We renormalize the field
$\phi\equiv\sqrt{Z}\phi_{r}.$ (12)
The resulting propagator for this renormalized field is
$D(p^{2})=\frac{i}{p^{2}-m^{2}(p^{2})},$ (13)
where we are naturally led to define the momentum-dependent mass [27], to ${\cal O}((p^{2}-m^{2})^{2})$
$m^{2}(p^{2})=m^{2}+\tilde{\Sigma}(p^{2})-\tilde{\Sigma}(m^{2})-(p^{2}-m^{2})\frac{d\tilde{\Sigma}}{dp^{2}}(m^{2}).$ (14)
As in the case of the quartic coupling (7), we expect the quantum correction to $m^{2}$ in (14) is small, at the leading order, we have formally free propagator
$D_{0}(p^{2})=\frac{i}{p^{2}-m^{2}}.$ (15)
It turns out to be tree-level or zero-loop order, so we put the subscript 0. Again, we may check that the combination $\tilde{\Sigma}(p^{2})-\tilde{\Sigma}(m^{2})-(p^{2}-m^{2})d\tilde{\Sigma}/dp^{2}(m^{2})$ is finite and parametrically small [3, 4, 6, 5].
The success of QFT with the renormalization is the prediction of the scale dependence of the physical parameter. We specify the scale by the external momentum $p^{2}$ that can be measured by the experiment. What we observe from the running gauge coupling is its dependence on the external momentum. If one insists, we may generalize the concept of the mass as that in (14), however, it is not separately observable. It is clear that only the whole combination in the propagator is physical, and it is not meaningful to mention a part of (13). We can measure energy-dependent mass $m^{2}(p^{2})$ from the reference $m^{2}$ without reference to the bare mass $m_{B}^{2}$ [3] by scattering experiments.
Figure 2: The (loop-order-)weighted Feynman diagram for the one-loop corrected propagator in $\phi^{4}$ theory. Although the two-point correlator receives the one-loop correction from the amplitude corresponding to the second diagram in the RHS, the propagator is modified as the geometric sum of its powers. Therefore, we regard the total sum as the one-loop correction to the propagator $D_{(n)}(p)^{2}$. The effective mass $m_{(1)}^{2}(p^{2})$ is defined as momentum-dependent, accordingly. Every subdiagram in this paper is renormalized and finite. In the $\phi^{4}$ theory, all the corrections vanish.
We calculate the self-energy perturbatively. The one-loop corrected propagator diagram is drawn in Fig. 2. The one-loop correction is
$\begin{split}-i\tilde{\Sigma}_{1}(p^{2})&=\frac{1}{2}\int\frac{d^{4}k}{(2\pi)^{4}}(-6i\lambda)D_{0}(k^{2})\\ &=\int\frac{d^{4}k}{(2\pi)^{4}}\frac{3\lambda}{k^{2}-m^{2}}.\end{split}$ (16)
It is independent of the external momentum. Thus, it is renormalized to zero. It is a well-known speciality in the $\phi^{4}$ theory and spontaneously broken non-Abelian gauge theories. Otherwise, we may have this loop inside the general loop, which cannot be removed by hand [6]. The nontrivial contributions arise from the two-loop order that we calculate below.
2.2 Two-loop propagator from effective couplings
Figure 3: Weighted Feynman diagram for two-loop corrected propagator in $\phi^{4}$ theory, having the momentum-dependent mass and its expansion. We denote the resulting propagator like one on the LHS, drawing a line and putting the loop correction order below. Two-loop corrections are obtained by replacing the quartic coupling with the physical one-loop corrected one $\lambda_{(1)},m^{2}_{(1)}$. The one-loop corrected quartic coupling $\lambda_{(1)}$ receives contributions from $s,t,u$-channels. The effect of the one-loop corrected propagator is on the second line, and that of the one-loop corrected vertex is on the third line. Note that there can be identical diagrams from different origins.
A loop subamplitude plays the role of the effective coupling, which is running. In this section, we show that one-loop corrected effective couplings enable us to write down two-loop amplitude as a “weighted one-loop” amplitude.
Since we have regularization-independent physical quartic coupling, it has well-defined loop corrections
$\lambda({\underline{p}})=\sum_{l=0}^{\infty}\lambda_{l}({\underline{p}}),$ (17)
where $\lambda_{l}(p^{2})$ is the $l$-loop correction and we define $\lambda_{0}(p^{2})\equiv\lambda.$ In place of the tree-level coupling $\lambda$, we may use one-loop correction $\lambda_{1}(p^{2})$ in (7). We can formally write down a “one-loop” amplitude using this
$-i\tilde{\Sigma}_{2}(p^{2})=\frac{1}{2}\int\frac{d^{4}k}{(2\pi)^{4}}[-6i\lambda_{1}(p,-p,-k,k)]\frac{i}{k^{2}-m^{2}}.$ (18)
Letting $p$ be the external momentum, the momenta $(p,-p,-k,k)$ flow in the vertex, summing to zero. From (7), the coupling $\lambda_{1}$ has three channels $s=(p+k)^{2},t=0,u=(p-k)^{2}$. Thus the equation (18) means that
$\tilde{\Sigma}_{2}(p^{2})=\frac{(6\lambda)^{2}}{2}\int\frac{d^{4}k}{(2\pi)^{4}}[V(s)-V(s_{0})+V(t)-V(t_{0})+V(u)-V(u_{0})]D_{0}(k^{2}).$ (19)
This is the sum of amplitudes at two-loop order without subdiagram divergences. They are all obtained by using the one-loop renormalized quartic coupling (and mass) (see also [25]). The only UV divergence arises from the large $k^{2}$ integral (not the $V(s)-V(s_{0})$, etc.), and it is the overall divergence of the “one-loop” amplitude that we remove shortly.
The $s$- and $u$-channel amplitudes in (19) are the famous “sunset” diagram that are renormalized.. In each, a potential “subdiagram” divergence in the first term is canceled by the second term, regarded as the counterterm contribution. The $t$-channel contribution is zero because $V(t)=V(t_{0})$ is independent of the external amplitude.
The sunset diagram is famous because of the overlapping internal lines shared by different loops. Our construction has two interesting points. One has good and bad sides. The corresponding diagram has “enhanced” symmetry, that is, the shuffling among three internal loops. This is unexpected because the one-loop diagram, which locally replaces the constant coupling, does not see the global structure of the resulting diagram. The sunset diagram has $3!$ permutation symmetry among the loops. So, we modify the symmetry factor by multiplying $1/3!$.
However, the other side of this is, in fact, the merit of this method. It originates exactly from the symmetry enhancement, and we see where the divergences arise: the one-loop structure of the integral (20) and the effective coupling $\lambda_{1}(p,-p,-k,k)$. We have no further divergence from overlapping effects. The use of effective coupling makes the symmetric diagram asymmetric in an intelligible way.
It does not have the flipping symmetry the original one-loop diagram has. Another interesting point is that this is expected and automatically taken into account in the full amplitude. The $u$-channel diagram is the flipped diagram but has the identical amplitude so that it correctly cancels the symmetry factor $1/2$
$\begin{split}\tilde{\Sigma}_{2}(p^{2})&=\frac{(6\lambda)^{2}}{2\cdot 3!}\int\frac{d^{4}k}{(2\pi)^{4}}[V((p+k)^{2})-V(s_{0})+V((p-k)^{2})-V(u_{0})]D_{0}(k^{2})\\ &=\frac{(6\lambda)^{2}}{3!}\int\frac{d^{4}k}{(2\pi)^{4}}[V((p+k)^{2})-\frac{1}{2}V(s_{0})-\frac{1}{2}V(u_{0})]D_{0}(k^{2}).\end{split}$ (20)
The constant term will be cancelled later, replacing $V(u_{0})$ with $V(s_{0})$ gives the same answer.
Naturally, the loop structure induces nontrivial momentum dependence in $\lambda_{(1)}$, so we have more than one correction amplitude. However, the total amplitude, being Lorentz scalar in $p$, only depends on the size of the external momentum $p^{2}$ after performing the integration over $k$. We know this using the symmetry $k\leftrightarrow-k$, it is an even function in $p$, $\tilde{\Sigma}_{(2)}(-p)=\tilde{\Sigma}_{(2)}(p)$ and depends on the even powers of $p$. Indeed, we have
[…]
with a $p^{2}$-dependent function
$F_{k^{2}}(p^{2})=\left(m^{2}-x(1-x)(p^{2}-k^{2})\right)^{2}-4x^{2}(1-x)^{2}m^{2}k^{2}.$
Using this modified vertex means that, in the Feynman diagram, we replace the vertex $\lambda_{(1)}({\underline{p}},k)$ with the one-loop corrected diagram as the last three diagrams in Fig. 3.
We may consider similar contributions from the propagator correction
$\frac{1}{2}\int\frac{d^{4}k}{(2\pi)^{4}}[-6i\lambda_{0}(p,-p,-k,k)]\frac{i}{k^{2}-m_{(1)}^{2}(k^{2})},$ (22)
which gives rise to further loop diagrams in the second line in Fig. 3. However, they do not depend on the external momenta and thus are all zero.
The final renormalized mass up to two-loop order is given in the renormalization condition (14) with $n=2$,
$\begin{split}m^{2}_{(2)}(p^{2})&=m^{2}+\tilde{\Sigma}_{(2)}(p^{2})-\tilde{\Sigma}_{(2)}(m^{2})-(p^{2}-m^{2})\frac{d\tilde{\Sigma}_{(2)}}{dp^{2}}(m^{2})\\ &=\sum_{l=0}^{2}m_{l}^{2}(p^{2}),\end{split}$ (23)
where
$m_{0}^{2}(p^{2})$ $=m^{2},$ (24)
$m_{1}^{2}(p^{2})$ $=0,$ (25)
[…]
Note that the dependence on $s_{0}$ and $u_{0}$ has disappeared, as desired; in fact, the coupling $\lambda$ is matched by the one-loop relation (6) at this point. We also see that in our scheme the mass $p^{2}=m^{2}$ is physical and unchanged $m_{(2)}^{2}(m^{2})=m^{2}$. However, the effective mass has a nontrivial profile in $p^{2}$ away from it. Accordingly, we can define a similar propagator $D_{(2)}(p^{2})$ up to two loops.
2.3 Two-loop quartic coupling from effective couplings
Figure 4: Two-loop corrections for the quartic coupling in $\phi^{4}$ theory using weighted one-loop Feynman diagrams. We obtain them by replacing one of the vertices in a one-loop diagram ($s,t,u$ respectively, rows) with the one-loop renormalized $\lambda_{1}({\underline{p}})$ $s,t,u$-vertices (columns). Replacing the other vertices gives similar diagrams, which are line-reflected diagrams of these. All the subdiagram parts corresponding to $\lambda_{1}({\underline{p}})$ are renormalized and finite.
In the same way, we calculate the two-loop corrections to the quartic coupling. We use one of the quartic couplings with the one-loop modified one $-6i\lambda_{(1)}({\underline{p}},k)$ in (7),
$\begin{split}(-6i\lambda)^{3}iV_{2}({\underline{p}})&=\frac{1}{2}\int\frac{d^{4}k}{(2\pi)^{4}}[-6i\lambda_{1}({\underline{p}},k)]D_{0}((p+k)^{2})D_{0}(k^{2})(-6i\lambda)\\ &\ +\frac{1}{2}\int\frac{d^{4}k}{(2\pi)^{4}}[-6i\lambda]D_{0}((p+k)^{2})D_{0}(k^{2})(-6i\lambda_{1}({\underline{p}},k)).\end{split}$ (27)
Here and hereafter, we denote the set of external momenta as ${\underline{p}}$ without confusion. Since all the factors in the integrand commute, the two terms give the same result. As before, there is no one-loop correction to the scalar mass.
Each vertex lies inside the loop with the internal momentum $k$, to which four momenta $({\underline{p}},k)\equiv\{-k,-p_{i},-p_{j},k+p_{i}+p_{j}\}$ flow in, preserving the total momentum. Here, $p_{i}$ and $p_{j}$ $(i\neq j)$ originate from the channel of the original one-loop diagram as in (4). Each vertex also has three independent channels from ${\underline{p}}$. Therefore, we have three possible channels in the subdiagram as functions of $s^{\prime}=(p_{i}+k)^{2},t^{\prime}=(p_{j}+k)^{2},u^{\prime}=(p_{i}+p_{j})^{2}$, respectively:
$\begin{split}V_{2}(s)&=\frac{1}{2}\int\frac{d^{4}k}{(2\pi)^{4}}\left(\left[V(s)-V(s_{0})\right]+\left[V((p_{4}+k)^{2})-V(t_{0})\right]\right.\\ &\qquad\left.+\left[V((p_{3}+k)^{2})-V(u_{0})\right]\right)D_{0}((p_{3}+p_{4}+k)^{2})D_{0}(k^{2}),\end{split}$ (28)
$\begin{split}V_{2}(t)&=\frac{1}{2}\int\frac{d^{4}k}{(2\pi)^{4}}\left(\left[V((p_{1}+k)^{2})-V(s_{0})\right]+\left[V(t)-V(t_{0})\right]\right.\\ &\qquad\left.+\left[V((p_{3}+k)^{2})-V(u_{0})\right]\right)D_{0}((p_{1}+p_{3}+k)^{2})D_{0}(k^{2}),\end{split}$ (29)
$\begin{split}V_{2}(u)&=\frac{1}{2}\int\frac{d^{4}k}{(2\pi)^{4}}\left(\left[V((p_{1}+k)^{2})-V(s_{0})\right]+\left[V((p_{4}+k)^{2})-V(t_{0})\right]\right.\\ &\qquad\left.+\left[V(u)-V(u_{0})\right]\right)D_{0}((p_{1}+p_{4}+k)^{2})D_{0}(k^{2}).\end{split}$ (30)
We have two-loop corrections as a result. The corresponding diagrams (accounting for $1/2$ of the corresponding amplitudes) are drawn in Fig. 3. Remarkably, all of these corrections are expressed by the single function $V_{2}(p^{2})$. They are arranged in terms of the Mandelstam variables, which are the only scalar quantities made of the conserved momentum. Indeed, in the diagram, we can verify that every diagram matches another in each row after changing the order of the external momenta. It is highly nontrivial from the asymmetric dependence on the external momenta to see that each integral is the function of Mandelstam variables. For instance, in (28), it is known that the finite parts of the second and the third bracketed terms are functions of $s$ only [26, 28]. Thus, they are the same. This cancels the symmetric factor that existed in the one-loop. Also, the first bracketed term in (28) has the correct symmetry factor $1/(2\cdot 2)$, from flipping two one-loop parts (one $1/2$-factor is inside $V(s)$).
We have obtaind the quartic couplings up to two-loop corrections
$\lambda_{(2)}({\underline{p}})=\lambda_{B}Z^{2}+\sum_{p^{2}=s,t,u}\big{[}6\lambda^{2}V(p^{2})+36\lambda^{3}V_{2}(p^{2})\big{]}.$ (31)
Again, there is no subdiagram divergence because all the sub-loops are contained in the modified coupling $\lambda_{1}({\underline{p}},k)$ are renormalized. The renormalization condition for the total amplitude is the same as before in (5),
$\begin{split}\lambda&=\lambda_{(2)}\left(s_{0},t_{0},u_{0}\right)=\lambda_{B}Z^{2}+\sum_{p_{0}^{2}=s_{0},t_{0},u_{0}}\big{[}6\lambda^{2}V(p_{0}^{2})+36\lambda^{3}V_{2}(p_{0}^{2})\big{]}+{\cal O}(\hbar^{3}\lambda^{4}).\end{split}$ (32)
Note that we need to match again the bare coupling $\lambda$, valid at ${\cal O}(\hbar^{2}\lambda^{3})$, instead of putting the relation (6), which is not accurate enough to this order. It is because in calculating $V(p^{2})$ and $V_{2}(p^{2})$, which also depend on $\lambda$, we have use more precise coupling $\lambda$ satisfying the equation (32), which is different from that in (6) (otherwise the order-2 correction should be zero). Thus the physical coupling at two-loop order is
$\lambda_{(2)}({\underline{p}})=\sum_{l=1}^{2}\lambda_{l}({\underline{p}}),$ (33)
where
$6\lambda_{0}({\underline{p}})$ $=6\lambda,$ (34)
$6\lambda_{1}({\underline{p}})$ $=\hbar(6\lambda)^{2}\bigg{[}\sum_{p^{2}=s,t,u}V(p^{2})-\sum_{p_{0}^{2}=s_{0},t_{0},u_{0}}V(p_{0}^{2})\bigg{]},$ (35)
$6\lambda_{2}({\underline{p}})$ $=\hbar^{2}(6\lambda)^{3}\bigg{[}\sum_{p^{2}=s,t,u}V_{2}(p^{2})-\sum_{p_{0}^{2}=s_{0},t_{0},u_{0}}V_{2}(p_{0}^{2})\bigg{]}.$ (36)
Again, the bare parameter $\lambda_{B}$ has disappeared and the additional loop correction at ${\cal O}(\hbar^{2}\lambda^{3})$ is given by the difference of the quantity $V_{2}({\underline{p}})$ at different scales, removing the dependence on the cutoff.
This shows the pattern of loop corrections for the quartic coupling. As we consider higher order terms, the renormalized higher correction adds, which we expect from the perturbative expansion. As a bonus, we learn some symmetry, for instance, that $V_{(L+1)}(s)$ as in Eq. (28) is a function on $s$ and $\lambda_{(L)}$ has shuffling symmetry under $s,t,u$, which we cannot easily know from the brute-force generation. By induction, as the complete sum, a physical parameter has the same symmetry.
2.4 Weighted Feynman diagram
We have seen that the two-loop corrected amplitudes are obtained from one-loop diagrams regarding the renormalized loop subdiagram as the effective coupling. This self-similar structure has become possible because the looped subdiagram is renormalizable and made finite.
For the Feynman diagram, we define a subdiagram $\gamma$ of a given diagram $\Gamma$ as a nonempty diagram whose sets of the vertices and lines are, respectively, subgroups of the ones in $\Gamma$. An effective mass gives a propagator as a subdiagram.
Starting from a two-loop diagram $\Gamma$, we may “contract” a one-loop subdiagram $\gamma$ and replace it with a vertex $\Gamma/\gamma$, symbolizing the generalized momentum-dependent coupling. That is, we regard the corresponding amplitude as formally a one-loop diagram with the “loop-order one” coupling $\lambda_{1}({\underline{p}},k)$. This gives us weighted Feynman diagram; we denote the orders in the vicinity of the corresponding vertex or lines. [^2] Without weights, it is known as a reduced Feynman diagram [17]. We may reflect the self-similarity of the loop-corrected couplings in the Feynman diagrams. Examples are given in Figs 2, 3, 4 and more to come.
Note that, to do this, all the subdiagrams have to be renormalized. As we have done in this section, if we perform contraction from the smallest loop diagrams to larger and renormalize them iteratively, we expect to have finite amplitudes.
3 Self-similar structure
In the previous section, we have generated two-loop couplings $\lambda_{2}({\underline{p}})$ and $m^{2}_{2}({\underline{p}})$ as the “one-loop” amplitude by replacing the coupling to one-loop amplitude $\lambda_{1}({\underline{p}})$. Such replacement is best defined in terms of the effective couplings and the coefficients of the effective action.
3.1 Effective couplings
The renormalized couplings and propagators are best understood as the effective couplings. We shall introduce an effective coupling as a vertex function in the momentum space that is renormalized. Following the standard functional method [30], we have the generating functional through the path integral
$Z[J]=e^{-iE[J]}=\frac{1}{Z[0]}\int{\cal D}\phi\exp\left[\frac{i}{\hbar}\int d^{4}x\left({\cal L}[\phi]+J\phi\right)\right].$ (37)
Here, ${\cal L}[\phi]$ is the Lagrangian defining the theory. Evaluating the path integral over $\phi$ gives the free energy $E[J]$ as an effective action containing the quantum corrections. Defining the classical field $\overline{\phi}$ and its fluctuation $\hat{\phi}$
$\overline{\phi}(x)\equiv-\frac{\delta E[J]}{\delta J(x)},\quad\phi(x)\equiv\overline{\phi}(x)+\hat{\phi}(x),$ (38)
the Legendre transform replaces the dependent variable $J$ with $\overline{\phi}$, yielding the usual effective action
$\Gamma[\overline{\phi}]=-E[J]-\int d^{4}xJ(x)\overline{\phi}(x),\quad\quad\frac{\delta\Gamma[\overline{\phi}]}{\delta\overline{\phi}}=-J.$ (39)
The effective action is expanded in the classical fields $\overline{\phi}(x)$, which defines $n$-point vertex functions
$\Gamma^{(n)}(x_{1},\dots,x_{n})=\frac{\delta^{n}\Gamma[\overline{\phi}]}{\delta\overline{\phi}(x_{1})\dots\delta\overline{\phi}(x_{n})}\bigg{|}_{J=0}.$ (40)
In this paper, we consider renormalizable theory, which means that the dimension of each term is less than or equal to four. Nevertheless, we may have high-order effective interaction $n>4$.
It is well-known that the $n$-point effective action is the same as the 1-particle irreducible $n$-point correlation function
$\Gamma^{(n)}(x_{1},\dots,x_{n})=-i\langle\phi(x_{1})\dots\phi(x_{n})\rangle_{\text{1PI}}$ (41)
that completely expands the theory [26]. Also, this is the minimal unit that appears in the $S$-matrix. (In fact, we need further field-strength renormalization for quantities with $\delta>0$, see below.) Therefore, we can use the effective action that defines the theory.
We use the coupling in the momentum space
$(2\pi)^{4}\delta^{(4)}(p_{1}+\dots+p_{n})\Gamma^{(n)}({\underline{p}})=\int\frac{d^{4}p_{1}}{(2\pi)^{4}}\dots\frac{d^{4}p_{n}}{(2\pi)^{4}}\Gamma^{(n)}(x_{1},\dots,x_{n})e^{-i(p_{1}\cdot x_{1}+\dots+p_{n}\cdot x_{n})}.$ (42)
Except for the bare coupling, the coupling is expanded in loop order
$\Gamma^{(n)}({\underline{p}})=\Gamma^{(n)}_{0}+\sum_{l=1}^{\infty}\Gamma^{(n)}_{l}({\underline{p}}),\quad\Gamma^{(n)}_{l}({\underline{p}})\propto\hbar^{l},$ (43)
where $\Gamma^{(n)}_{l}({\underline{p}})$ is the $l$-loop correction. In particular, the zeroth order coupling is the bare parameter in the Lagrangian
$\Gamma^{(n)}_{0}=g^{(n)}_{B}$ (44)
In the $\phi^{4}$-theory, we also have $\Gamma^{(n)}_{l}({\underline{p}})\propto\lambda^{l}$. An obvious but important property shared by all order couplings is momentum conservation
$p_{1}+\dots+p_{n}=0,\quad\text{ for all $l$}.$ (45)
We calculate the couplings using the standard perturbation order by order (see e.g. [26]). Separating the action to the free $S_{0}$ and the rest $S_{\text{int}}$ and taking $Z_{0}[J]=\int{\cal D}\phi e^{-S_{0}[\phi]}$, we construct
$Z[J]=\exp\left(-S_{\text{int}}\left(\frac{\delta}{\delta J}\right)\right)Z_{0}[J].$ (46)
At order $L$, we construct the effective coupling by applying the Feynman rules, see (59) and Section 3.3 below. The corresponding Feynman diagram is characterized by $n$ external scalar field. In general, there is more than one diagram contributing to $\Gamma_{L}^{(n)}({\underline{p}})$, as we have seen in Fig. 4 for $\Gamma_{2}^{(4)}(p_{1},p_{2},p_{3},p_{4})$, for example. We can also generalize it to a theory with fields of various spins, in which case the superscript should contain more information on the external fields.
3.2 The problem of full renormalization
Now, we wish to compare the theory with the experiment. We fix the physical coupling at a reference scale $p^{2}=\mu^{2}$ by comparing it with the matrix element at ${\underline{p}}={\underline{\mu}}$
$-i\Gamma^{(n)}_{(L)}(\underline{\mu})=i{\cal M}(\text{$n$ scalars})=-i\Gamma^{(n)}_{\text{phy}}+{\cal O}(\hbar^{L+1}),$ (47)
generalizing the quartic coupling in (5). Note that to physical parameter $\Gamma^{(n)}_{\text{phy}}$ in fact depends on the accumulated order $(L)$; Adding higher order correction predicts the more accurate theoretical value $\Gamma^{(n)}_{(L)}(\underline{\mu})$ fixing $\Gamma^{(n)}_{\text{phy}}$ more precisely.
The premise that we can take the renormalization condition (47) may be a big assumption. We should guarantee that unknown UV physics beyond our reach must not affect the condition. The Appelquist–Carazzone decoupling theorem [29] and its extension to the scalar mass [3, 4] justify it.
Our original theory can be alternatively parameterized using this physical parameter. We can express the parameters using the reference one, eliminating the bare one
$\Gamma^{(n)}({\underline{p}})=\Gamma^{(n)}_{\text{phy}}+\sum_{l=1}^{\infty}\left(\Gamma^{(n)}_{l}({\underline{p}})-\Gamma^{(n)}_{l}({\underline{\mu}})\right).$ (48)
In Section 2, we took the reference scale $\mu^{2}=s_{0},t_{0},u_{0}$ at which the coupling $\lambda$ is defined. For the mass, we took the reference scale $\mu^{2}=m^{2}$ as the pole mass. Since we expand the theory around these, the scales $s_{0},t_{0},u_{0},m^{2}$ should not be different too much.
Although the leading divergence in $\Gamma^{(n)}_{l}({\underline{p}})$ is removed in this expression, we still have the divergence. This is measured by the superficial degree of divergence $\delta(\gamma)$ of the amplitude associated with a diagram $\gamma$ is defined as [31]
$\delta(\gamma)=4-\sum_{f}(s_{f}+1)E_{f}+\sum_{i}N_{i}\Delta_{i},$ (49)
where $f$ labels the external legs and $i$ labels the vertices so that it does not depend on the details of the internal part of the Feynman diagram. Also, $\Delta_{i}$ is the mass dimension of the corresponding physical parameter, and $E_{f},s_{f}$ are the numbers of external legs and the spin of the field $f$, respectively. In what follows, we may omit the $\gamma$-dependence and simply denote $\delta$ without confusion.
Also, we have seen that for the loop corrections to the quartic coupling, it was $\delta=0$, and just the subtraction worked. However, for the scalar mass, the subtraction was not enough to remove the subleading divergence. Understood as the order-$\delta$ polynomial, in general, every coefficient of $\Gamma^{(n)}_{l}({\underline{p}})$ is divergent. This is the content of the Weinberg theorem.
Theorem 1
(Weinberg [25]) The degree $\delta$ is the degree of the divergence of the momentum integral if all the momenta go to infinity. Thus, the amplitude is UV finite if $\delta<0$.
That is, the corresponding amplitude behaves like $k^{\delta}$ in the large momentum limit. It is nontrivial to separate the large momenta of the subdiagrams and those of the main diagram. The theorem shows that the large momentum divergence of the overall amplitude is determined by the external legs of the diagram.
An amplitude formed by renormalizable interactions and involving a loop is divergent if the number of external legs is equal to or smaller than the number of dimensions, 4. The divergent diagram can be absorbed into the coupling. This justifies the loop-corrected diagrams for renormalizable ones.
We see that the condition we impose is the same analytic behavior as the free particle. This is to impose the self-similarity for the loop-corrected scalar field to behave as if free particle. It is realized by Taylor expansion and comparison. Let $\delta$ be the degree of divergence of the amplitude $\Gamma^{(n)}_{L}$. Using $\Gamma^{(n)}_{L}({\underline{p}})$, we have the renormalized coupling [33, 17]
$\overline{\Gamma}^{(n)}_{(L)}({\underline{p}})=\Gamma^{(n)}_{\text{phy}}+\left[1-t^{\delta}_{\underline{\mu}}\right]\Gamma^{(n)}_{L}({\underline{p}}),$ (50)
where $t^{\delta}_{\underline{\mu}}$ is the Taylor expansion operator with the order $\delta$ around the reference momentum ${\underline{p}}=\underline{\mu}$,
$t^{\delta}_{\underline{\mu}}\Gamma^{(n)}_{(L)}({\underline{p}})=\sum_{i=0}^{\delta}\frac{1}{d!}(p_{1}^{\alpha_{1}}-\mu_{1}^{\alpha_{1}})\dots(p_{i}^{\alpha_{i}}-\mu_{i}^{\alpha_{i}})\frac{\partial^{i}\Gamma^{(n)}_{(L)}}{\partial p_{1}^{\alpha_{1}}\dots\partial p_{i}^{\alpha_{i}}}({\underline{p}}=\underline{\mu},{\underline{k}}).$ (51)
We need $\delta\geq 0$ for renormalizability. The amplitude $\Gamma^{(n)}_{(L)}$ contains divergence if we differentiate $i<L$ times. To compare with the actual measured value $\Gamma^{(n)}_{\text{phy}}$, extracting finite quantity from $L$-loop amplitude is not enough, but we should also fix the physical value; We need the complete amplitude in $\Gamma^{(n)}_{(L)}$.
Our definition of the renormalized parameter can be transformed from the BPHZ parameter by the addition of physical mass and finite renormalization
$\left[1-t^{\delta}_{\underline{\mu}}\right]\Gamma^{(n)}_{(L)}=\left[1-t^{\delta}_{\underline{p}}\right]\Gamma^{(n)}_{(L)}+\left[t^{\delta}_{\underline{p}}-t^{\delta}_{\underline{\mu}}\right]\Gamma^{(n)}_{(L)}$ (52)
where all the (three pairs of) terms are finite, having the subtraction structure made finite by Taylor expansions.
The Taylor expansion structure originates from the field strength renormalization as in the derivation of the momentum-dependent mass (14),
$\phi\equiv\sqrt{Z}\phi.$ (53)
This is the redefinition of the field, which is, in the $\phi^{4}$-theory, the only field that should affect all the parts of the amplitude. What the BPHZ scheme does is use the strength reormalized field. After the renormalization, we have ${\cal L}[\phi]={\cal L_{\text{ren}}}[\phi_{r}],J\to\sqrt{Z}J$ and
$Z[J]=e^{-iE[J]}=\frac{1}{N^{\prime}}\int{\cal D}\phi_{r}\exp\left[\frac{i}{\hbar}\int d^{4}x\left({\cal L_{\text{ren}}}[\phi_{r}]+J\phi_{r}\right)\right],$ (54)
with the normalization $N^{\prime}$ obtained by setting the new current to zero. We use the strength renormalized field $\phi_{r}$, and the rest of the perturbation is the same.
Although the field-strength renormalization procedure (53) should remove all the divergence by counterterms, we do not know such counterterms before calculating the amplitude. In practice, we use the effective action before field-strength renormalization (40) and obtain counterterms using the renormalization condition. Then, we can remove the overall divergence using the procedure (50). Since this procedure alone cannot remove the subdiagram divergence, we have to apply it to subamplitudes.
Traditionally, the last terms in (50) or the Taylor expanded part have been called counterterms. The counterterm interactions should be regarded as a part of the bare Lagrangian, not additionally introduced ones when necessary; then, we have uncancelled terms. After loop corrections and renormalization, we have the momentum-dependent terms. It is important to note that, due to the subtraction structure (50), the whole effective action, and hence the couplings $\overline{\Gamma}_{L}({\underline{p}})$ are finite, independent of regularization [17], and parametrically small [3, 4]. Unrenormalized amplitudes $\Gamma_{L}^{(n)}$ do not have this feature, so they cannot play the role of effective coupling in perturbation.
Higher order terms in $p_{i}$’s are given by Taylor expansion coefficients, fixing, for instance, the normalization of the fields that we cannot observe, in view of the examples in (14) and (23). Technically, although the operator $t^{\delta}$ removes the divergence, physically, it means that the behavior of physics should be unchanged in the vicinity of the renormalization point ${\underline{p}}=\underline{\mu}$ to ${\cal O}(|{\underline{p}}-\underline{\mu}|^{n}$). This procedure matches and removes the bare coupling up to ${\cal O}(\hbar^{L+1})$, as well as the divergent part.
Having renormalized the main part, the remaining problem is that subamplitudes in $I^{(n)}_{L}$ may contain divergences. Dyson suggested renormalizing the subamplitude recursively [33]. After renormalizing the main amplitude, we go down to check the subdiagram divergence. If present in $\gamma$, we renormalize the corresponding (sub)amplitude. Eventually, this $\gamma$ may contain subdiagram divergence again, but repeating the replacements will make it finite, and the procedure ceases after finite steps.
The program becomes procedural thanks to the Bogoliubov relation, which is the general recursive procedure for removing the subdiagram divergence [14].
Theorem 2
(Bogoliubov and Parasuik [14], Hepp [15]) A given amplitude $A_{\Gamma}$ corresponding to a graph $\Gamma$ is renormalized by
$\overline{A}_{\Gamma}=(1-t^{\Gamma})A^{\prime}_{\Gamma},$ (55)
where $A^{\prime}$ is recursively defined as
$A^{\prime}_{\gamma}=A_{\gamma}+\sum_{\{\gamma_{1},\dots\gamma_{c}\}}A_{\gamma/\{\gamma_{1}\dots\gamma_{c}\}}\prod_{i=1}^{c}(-t^{\gamma_{i}})A_{\gamma_{i}},$ (56)
and $\{\gamma_{1},\dots,\gamma_{c}\}$ is a set of mutually disjoint connected subdiagrams of $\gamma$.
Eventually, Zimmermann’s forest formula [17] provides an explicit solution satisfying the recursive relation (56) and has become the canonical method in this top-down approach so far.
In this paper, we take the opposite, bottom-up route, starting from the smallest subloop. The main goal is to generate and completely renormalize effective couplings $\overline{\Gamma}^{(n)}_{(L)}({\underline{p}})$ with expansion
$\overline{\Gamma}^{(n)}_{(L)}({\underline{p}})=\overline{\Gamma}^{(n)}_{\text{phy}}+\sum_{l=1}^{L}\overline{\Gamma}^{(n)}_{l}({\underline{p}}),$ (57)
where $\overline{\Gamma}^{(n)}_{l}({\underline{p}})$ is the $l$-loop correction. We try to find the completely renormalized coupling iteratively.
Definition 1
We call the completely renormalized parameter $\overline{\Gamma}^{(n)}_{L}({\underline{p}})$ an effective parameter of (loop) order $L$. Here $\Gamma_{\text{phy}}^{(n)}$ is the value at a reference scale. We define also weight $L$ propagator using the renormalized mass $m_{(L)}({\underline{p}})$ inside the propagator
$\overline{D}_{(L)}(p^{2})=\frac{i}{\overline{\Gamma}^{(2)}_{(L)}({\underline{p}})},$ (58)
which generalizes the scalar mass (23).
Every other coupling $\overline{\Gamma}^{(n)}_{l}({\underline{p}})$ has the same structure of the momentum conservation so that they can play the role of the coupling.
The traditional approach to the complete renormalization program is bottom-up. Starting from the overall-renormalized amplitude, we track subamplitudes having divergence and apply the renormalization procedure.
In this paper, we take an alternative, bottom-up approach. We construct the amplitude with renormalized subamplitude. In both approaches, the self-similarity of loop amplitude is importantly used. A subdiagram is no different from the main amplitude in the sense that it itself and further sub-sub amplitude may have the same structure.
In the end, the only quantities we can observe via the $S$-matrix are these fully renormalized couplings $\overline{\Gamma}^{(n)}({\underline{p}})$. For instance, neither $m^{2}_{B},\tilde{\Sigma}(p^{2})$ or even the mass (14) is separately observable [3]. In fact, the order two coupling $\overline{\Gamma}^{(2)}(p^{2})=p^{2}+m^{2}(p^{2})$ or its inverse is the minimal unit appearing in the $S$-matrix.
For instance, the Higgs mass correction as a function of the external momentum can be compared with the experimental data [3]. The act of renormalization not only removes the divergence but also imposes the momentum dependence on the coupling [6].
3.3 Contraction of loops
With the effective couplings defined in the last subsection, we can generate fully renormalized amplitudes of an arbitrary loop order by iteration. Basically, we replace a coupling with a loop amplitude. To facilitate this, it is convenient first to define the reverse operation, the contraction, as replacing the loop amplitude with the constant coupling.
We have a Feynman diagram $\Gamma$, whose vertices $\{A\}$ and lines $\{l_{AB}\}$ correspond to the coupling and propagators, respectively, according to the Feynman rules. The closedness implies the set of serially connected lines, thus, a loop typically has many vertices along the lines of the loop. By a subdiagram $\gamma$ of a diagram $\Gamma$, we mean a closed loop, with vertices $\{a\}$ and lines $\{l_{ab}\}$ inside. Specifying a subdiagram as a set of lines is convenient because the vertices are automatically selected, while the opposite is not true. If a line connects two vertices of the whole diagram, it also connects those of the subdiagram.
Figure 5: A typical multiloop diagram. We can $r$-contract a subgraph $\gamma$ to a point $v$ as long as the resulting tree coupling exists in the Lagrangian. The total inflowing momenta $\{\ell^{v}_{i}\}$ to $\gamma$ and the same to $v$ is conserved.
A general amplitude, or one component of the unrenormalized effective coupling $\Gamma_{L}^{(m)}({\underline{p}})$, has the form [17]
$A^{(m)}_{L}({\underline{p}})=\int\frac{d^{4}k_{1}}{(2\pi)^{4}}\dots\frac{d^{4}k_{L}}{(2\pi)^{4}}\prod_{\{A\}}\big{[}g_{A}P_{A}({\underline{p}},\underline{\tilde{k}_{i}})\big{]}\prod_{\{l_{AB}\}}\Delta_{AB}({\underline{p}},\underline{k_{i}}).$ (59)
Here $\Delta_{AB}({\underline{p}},\underline{k_{i}})$ are propagators along the lines $l_{AB}$. At the vertices $A$, we also have constant couplings $g_{A}$ and polynomials $P_{A}({\underline{p}},\underline{\tilde{k}_{i}})$. [^3] We understand that the product reaches only the next factor, not all the rest factor on the right. They are physical ones in (47) used in the expansion.
At each vertex, there are inflowing external momenta $\ell_{i}^{A}$ to $v_{A}$. They are linear combinations of the external momenta $\{p_{i}\}$ and the subset $\{\tilde{k}_{i}\}\subset\{k_{i}\}$. [^4] There is the canonical way to relate the momenta, called substitution operation [17]; however, we do not need it. Its sum is conserved, like in (44),
$\sum_{i=1}^{n}\ell_{i}^{A}=0.$ (60)
We can (partly) perform the integrals in (59) to obtain functions. See Fig. 5. We partition the set of momenta into those of contracted graph $\gamma$$\{k^{\prime\prime}_{i}\}$ and the rest $\{k_{i}\}$
$\{k_{1},k_{2},\dots,k_{L}\}=\{k_{1}^{\prime},k_{2}^{\prime},\dots k_{L-v}^{\prime}\,|\,k^{\prime\prime}_{1}\dots k^{\prime\prime}_{v}\},$
with appropriate reordering and renaming. The sub-amplitude is separated such that
$A^{(m)}_{L}({\underline{p}})=\int\frac{d^{4}k^{\prime}_{1}}{(2\pi)^{4}}\dots\frac{d^{4}k^{\prime}_{L-v}}{(2\pi)^{4}}A_{v}^{(n)}({\underline{p}},\underline{k^{\prime}_{i}})\prod_{\{A\}-\{a\}}\big{[}gP({\underline{p}},\underline{\tilde{k}^{\prime}_{i}})\big{]}\prod_{\{l_{AB}\}-\{l_{ab}\}}\Delta({\underline{p}},\underline{k^{\prime}_{i}}),$ (61)
with
$A_{v}^{(n)}({\underline{p}},\underline{\tilde{k}^{\prime}_{i}})=\int\frac{d^{4}k^{\prime\prime}_{1}}{(2\pi)^{4}}\dots\frac{d^{4}k^{\prime\prime}_{v}}{(2\pi)^{4}}\prod_{\{a\}}\big{[}g_{a}P_{a}({\underline{p}},\underline{\tilde{k}^{\prime}_{i}})\big{]}\prod_{\{l_{ab}\}}\Delta({\underline{p}},\underline{k^{\prime\prime}_{i}}),$ (62)
such that there is no $k^{\prime\prime}_{i}$ dependent functions in the integrand in (61) any more. Then $A_{v}^{(n)}({\underline{p}},\underline{k^{\prime}_{i}})$ is a function of the external and the rest of the momenta.
The subamplitude depends on the inflowing external momenta $\ell_{i}^{a}$ attached to the loop $\gamma$ through the vertices $v_{a}$, which are linear combinations of the external momenta $\{p_{i}\}$ and the subset $\{\tilde{k}^{\prime}_{i}\}\subset\{k_{i}^{\prime}\}$. Again, its sum is conserved,
$\sum_{i=1}^{n}\ell^{a}_{i}=0.$ (63)
At every vertex, we have a zero-sum of the inflowing momenta. Especially when collecting the vertices attached to the external legs, the total momenta flowing into all the vertices vanishes because the total momenta inside the loop is zero. After the cancellation, we are left with the external momenta. Thus, we have (60) and (63). This should be, to make sense of (43). After the contraction, this conservation is interpreted as the momentum conservation on the vertex as (60) and (63).
From the structure of the effective couplings (57) on the same momentum conservation, we can exchange $A_{v}^{(n)}$ with any of $\Gamma_{j}^{(n)}$.
Definition 2
We define $r$-contraction as the replacement of the (sub)amplitude to a constant coupling
$A_{v}^{(n)}({\underline{p}})\to g_{B}^{(n)},\quad v>0,$ (64)
and the contraction is possible if we have a bare coupling $g_{B}^{(n)}$ in the defining Lagrangian in (37).
Clearly, the map is many to one. Ultimately, what we need is the reverse of the contraction, from a constant coupling $\Gamma_{0}^{(n)}=g_{B}^{(n)}$ to the loop amplitude like $A_{v}^{(n)}({\underline{p}},\underline{k}).$ Since we are going to calculate amplitude from the Lagrangian, we need a tree-level coupling.
In effect, we have the following theorem.
Theorem 3
The $r$-contractability is the same as the renormalizability of the overall amplitude if the Lagrangian is renormalizable and generic.
In renormalizable theory, we do not have a bare coupling for non-renormalizable operators. Thus, the contraction is limited to this case. However, it is different from renormalizability if the interaction Lagrangian is non-generic. By genericity, we mean the following.
Definition 3
A generic Lagrangian contains all the operators allowed by symmetry of the theory.
In the example of the theory with a massless scalar field, contraction may not be possible even for a renormalizable effective interaction.
In the graph side, the above $r$-contraction of a diagram $\gamma$ is to send all the vertices $\{v_{a}\}$ in $\gamma$ to one point $\{v\}.$ We disregard the lines $l_{ab}$ in $\gamma$. When contracting the loop in the propagator, we also drop the vertex. Considering amputated external lines of $\gamma$, the connectivity to $\{v_{i}\}$ to external lines is inherited. This graphical operation also allows the contraction of a line, but we limit our discussion to the contraction of a loop. $\gamma$ is not necessarily a single loop; a multi-loop can be contractable at once. A propagator that is not 1PI can also be contractable in the same way.
It can be that $\gamma$ is a subdiagram of a diagram $\Gamma$. We denote the contraction by
$\Gamma\to\Gamma/\gamma.$ (65)
This graph contraction does not distinguish $r$-contraction from a usual contraction, for which any of the loops with $n$-external lines are contractible, leaving a $n$-point vertex.
How do we formally distinguish between the main and subdiagrams? We introduce the notion of the reducibility of the $r$-contraction that we study in the next section.
3.4 Irreducible diagrams and non-renormalizable operators
Figure 6: Irreducible diagrams are characterized by an internal line connecting non-adjacent vertices. We can generate such ones (left) by flop change, namely, exchanging some vertices of diagrams of effective couplings (right). This process preserves the loop order.
In most cases, we can increase the loop order one by one. This dominates the generation of higher loop amplitudes. Consider an interaction of type
${\cal L}\supset XYY,\quad X=Y\text{ or }X\neq Y.$ (66)
This includes the self-coupling in the $\phi^{3}$-theory, the Yukawa coupling and the gauge couplings. It always forms a one-loop $n$-point amplitude of the field $X$, including the self-energy corrected by $X$. Also, the self-energy, as usual, and the quartic couplings from a box diagram can be made. Beyond (66), if an interaction has too many legs, it becomes difficult to obtain renormalizable interaction since, after taking one loop, we are left with many external legs. Also, if we have three fields with only interaction
${\cal L}\supset XYZ,\quad X\neq Y\neq Z\neq X,$ (67)
one-loop self-coupling can be made. However, order-3 or order-4 coupling is difficult to make, and we need extra couplings.
In non-generic cases, $r$-contraction of a one-loop subamplitude is not possible because it requires the existence of a corresponding constant coupling term in the Lagrangian, which may not exist. For instance, although we deal with renormalizable theory, we can form a one-loop amplitude with arbitrarily many external fields. For this graph, it is possible to shrink the loop to a point in a similar way as contraction, but the resulting operator becomes non-renormalizable. The corresponding amplitude is finite by default, but we do not have the corresponding coupling at the tree level.
Even if the total amplitude is renormalizable, a subdiagram is not always contractable. We call such an amplitude irreducible.
Definition 4
We define an irreducible amplitude as an $r$-contractable one in which none of the proper subamplitudes can be $r$-contracted. We define every $r$-contractable one-loop amplitude as irreducible. Otherwise, the amplitude is reducible.
Note that the irreducibility requires $r$-contractablity. The irreducible loop is the minimal amplitude requiring renormalization, having no subdiagram divergence from the construction.
Although the definition is about the amplitude, it is easier to consider the corresponding diagram. In the reducible multi-loop amplitude, the $r$-contraction of the subloop can be done. One example of the irreducible amplitude is a “Celtic-cross” amplitude in the $\phi^{4}$-theory, whose diagram is the last in Fig. 7, a four-loop diagram with four external legs. It cannot be obtained by expansion up to a three-loop ${\cal O}(\lambda^{3})$ weighted Feynman diagram because any loop subamplitude cannot be $r$-contracted. Another example is a three-loop amplitude with no crossing at one vertex of the Cletic-cross amplitude, which is on the first row in Fig. 7 [32]. They are the minimal irreducible amplitudes for the quartic coupling. Both have the overall logarithmic divergence $\delta=0$ and are free of subamplitude divergence.
An irreducible amplitude is characterized by subamplitudes with non-renormalizable effective couplings.
Theorem 4
Any subloop amplitude of irreducible amplitude is finite.
This follows from the definition of irreducibility and the Weinberg theorem. Any subamplitude $\gamma$ is non-renormalizable in renormalizable theory so that we have $\delta(\gamma)<0$.
4 Generating amplitudes using self-similarity
The above findings on the recursive structure of loop diagrams enable us to synthesize general loop amplitudes. We generalize the results in Section 2, where we have constructed all the quartic and quadratic couplings to two-loop order from one-loop amplitude by expansion.
4.1 Generating high loop-order couplings
We are ready to construct high-order amplitudes systematically. This can be done by some reverse process of $r$-contraction (64) accompanied by renormalization. An elementary unit of an additional loop is an irreducible amplitude defined in Definition 4. To it, the inverse of the contraction is well-defined as follows.
Definition 5
We define expansion of $g_{B}^{(n)}$ by $A_{\gamma}^{(n)}({\underline{p}})$ as the replacement of the tree-level coupling $g_{B}^{(n)}$ with the irreducible amplitude that is renormalized $\overline{A}_{\gamma}^{(n)}({\underline{p}})$
$g_{B}^{(n)}\to\overline{A}_{\gamma}^{(n)}({\underline{p}}).$ (68)
We do the same to the propagator.
For each coupling and each propagator, we have the expansion. Also, we may have more than one “channel” amplitude having the same external lines, specified by the irreducible diagrams $\gamma$.
On the graph side, an expansion at the vertex $v$ in a graph $\Gamma/\gamma$ by the graph $\gamma$ is the inverse of the contraction; we replace a $v$ with $\gamma$. The amplitude we have generated by expansion as above increases the loop order from zero to $l$, the loop order of $A_{\gamma}^{(n)}.$
Ultimately, we want a completely renormalized amplitude for a fixed loop order $L$. This can be done by summing the complete order-$L$ amplitudes. Since a number of different diagrams have the same contracted diagram, the contraction is many to one. For example, a loop order 10 diagrams can be made in various ways:
  • An order 7 diagram expanded by order one couplings at three vertices.
  • An order 6 diagram expanded by an order four irreducible coupling at one vertex.
  • An order 10 irreducible diagram.
  • and so on.
The processes do not commute because there are irreducible diagrams in various loop orders.
The total order of an amplitude related to a (sub)diagram $\gamma$ is invariant in the following sense.
Lemma 1
For a graph $\gamma$ having $m$-loop structure containing the order $l_{i}$ lines and order $v_{i}$ vertices. Then
$K=m+\sum l_{i}+\sum v_{i}.$ (69)
is an invariant after contraction followed by the expansion of the same order.
We may use this result for a total graph or a subgraph. We may call the above amplitude weighted $m$-loop graph.
By successive expansions on the couplings, we can scan the higher-order amplitudes. If we renormalize the overall amplitude whenever we expand, the resulting amplitude contains no subdiagram divergence. The only divergence comes from the total amplitude inherited from the amplitude before expansion. It can be readily renormalized. This guarantees divergence-free amplitudes at any loop order that we can compare with experiments.
Theorem 5
The following procedure, which we call generation, produces complete renormalizable amplitudes of the loop order $L$ that are fully renormalized,
$\overline{\Gamma}_{(L)}^{(n)}({\underline{p}}),\quad L>0.$ (70)
  1. 1.
    Start with an $n$-point coupling $\Gamma_{0}^{(n)}=g_{B}^{(n)}$. The mass $\Gamma_{0}^{(2)}$ determines a propagator $\Delta^{-1}_{0}(p^{2})$.
  2. 2.
    Assume we have all the renormalized $n$-point couplings $\overline{\Gamma}_{v}^{(n)}$ of order $v=1,\dots l-1$, for all $n$.
  3. 3.
    We obtain the order-$l$ reducible couplings as weighted $m$-loop amplitude
    $\begin{split}\Gamma^{(n),\text{r}}_{l}({\underline{p}})&=\sum_{m=1}^{l-1}\sum_{\text{channels}}\int\frac{d^{4}k_{1}}{(2\pi)^{4}}\cdots\frac{d^{4}k_{m}}{(2\pi)^{4}}\prod\left[\overline{\Gamma}^{(n_{i})}_{(l-m)}P_{(l-m)}\right]\prod\overline{\Delta}_{(l-m)}\bigg{|}_{l},\end{split}$ (71)
    from the $m$-loop integral couplings by the expansion of order-$(l-m)$ couplings, adopting the notation in (61). This is done by replacing the propagator $\Delta_{0}$ and the vertices $g^{(n)}_{B}$ with $\Delta_{(l-m)}$ and $\overline{\Gamma}_{(l-m)}^{(n)}$, respectively, and then extracting the order-$l$ contributions. The extraction can be done either by counting the number of loops in the corresponding diagrams or by counting the power of $\hbar$ in the amplitude.
  4. 4.
    Modify the symmetry factors from the symmetry of the resulting diagrams, if necessary. For each channel, we multiply the combinatoric factor
    (the order of reduced symmetry of the irreducible diagram used)(the order of enhanced symmetry after the $\frac{\text{(the order of reduced symmetry of the irreducible diagram used)}}{\text{(the order of enhanced symmetry after the expansion)}}.$ (72)
  5. 5.
    Collect all the irreducible amplitude at the order $l$,
    $\Gamma^{(n),\text{ir}}_{l}({\underline{p}})=\sum_{a}\Gamma^{(n),\text{ir},a}_{l}({\underline{p}}),$ (73)
    where the index $a$ labels the irregular amplitudes at the loop-order $l$. The channels and symmetry factor are as usual.
  6. 6.
    The total sum of the above amplitudes
    $\Gamma^{(n)}_{l}({\underline{p}})=\Gamma^{(n),\text{r}}_{l}({\underline{p}})+\Gamma^{(n),\text{ir}}_{l}({\underline{p}})$ (74)
    is the order-$l$ coupling having no subdiagram divergences.
  7. 7.
    Renormalize the overall amplitude in the same way as in Section 2.3, or using the formula (50),
    $\overline{\Gamma}_{(l)}^{(n)}({\underline{p}})=\Gamma_{\text{phy}}^{(n)}+\sum_{i=1}^{l-1}\overline{\Gamma}_{i}^{(n)}({\underline{p}})+\left[1-t^{\delta}_{\underline{\mu}^{(n)}}\right]\Gamma^{(n)}_{l}({\underline{p}}),$ (75)
    where $\delta$ is the degree of divergence originating of the amplitude $\Gamma_{l}^{(n)}.$ Note that, as in (57), the parenthesis in the subscript means the corresponding quantity is the accumulated quantities up to the denoted order. We use the same reference scale $\underline{\mu}^{(n)}$ for the same types of interactions $\overline{\Gamma}_{(l)}^{(n)}$. Note that the zeroth-order coupling depends on the order $l$ and thus needs to be re-matched.
  8. 8.
    Iterate from Procedure 2 to Procedure 7 until reaching the desired order $l=L$.
The method is inductive, and the generated diagrams are exhaustive because all other amplitudes than the reducible ones, inheriting the structure of lower-order amplitudes, are irreducible by definition. Since the irreducible diagram has no subloop divergence, it can be used as an elementary building block of expansion, increasing the loop order by one or more. At every stage, all the subdiagrams are divergence-free, and we always have the fully renormalized amplitude.
For the propagator, the geometric sum of the 1PI amplitudes $\tilde{\Sigma}(p^{2})$ gives the correction to the mass $m^{2}+\tilde{\Sigma}(p^{2})$. If we have a 1-particle reducible amplitude, we regard it as part of the geometric sum.
What if we have the following connection of two different 1-particle-reducible (1PR) amplitudes? Consider the amplitude of the schematic form
$D_{(l)}\tilde{\Sigma}_{a}D_{(l)}\tilde{\Sigma}_{b}D_{(l)},$ (76)
where $D_{(l)}$ is a propagator and $\tilde{\Sigma}_{i}$ is a self-energy. We may use an 1PI amplitude $\Sigma_{c}$ containing both $\tilde{\Sigma}_{a}$ and $\tilde{\Sigma}_{b}$. Its replacement
$D_{(l)}\tilde{\Sigma}_{c}D_{(l)}\tilde{\Sigma}_{c}D_{(l)}$ (77)
contains the original one (76), giving the correction
$m_{(l)}^{2}(p^{2})\to m_{(l)}^{2}(p^{2})+\tilde{\Sigma}_{c}(p^{2}).$ (78)
We may think of the “least common multiple” (LCM) amplitudes. One way to make the LCM amplitude is to replace one propagator of $\Sigma_{a}$ to contain $\Sigma_{b}$, or vice versa.
Theorem 5 is stated for a theory with a scalar. However, it is not difficult to generalize to one with more fields with different spins since the generic amplitude has a similar structure.
Summarizing, we have calculated the perturbative effective action up to order $L$ that is fully renormalized. Since it spans the Green’s function in QFT, we have the following theorem.
Theorem 6
Although the Lagrangian in QFT contains (bare) parameters related to infinity, no infinity appears in the renormalized effective action.
Figure 7: The complete quartic coupling at three-loop and four-loop. Most of the order three amplitudes are obtained by calculating one-loop integrals using order two couplings. There are, in general, irreducible diagrams at high loop order. Order four diagrams consist of the order-three diagrams with the vertices and lines modified plus irreducible ones. The square bracket is the extraction operation.
4.2 Examples
For instance, as in (27), the general order-$3$ quartic coupling is obtained as a weighted one-loop amplitude. Reducible amplitudes in the $s$-channel are
$\begin{split}V_{3}^{\text{r}}(p_{1},p_{2},p_{3},p_{4})&=\sum_{v_{1}+v_{2}+l_{1}+l_{2}=3}\frac{1}{2}\int\frac{d^{4}k}{(2\pi)^{4}}[-6i\lambda_{v_{1}}(p_{1},p_{2},k,-k-p_{1}-p_{2})]\\ &\quad\times[-6i\lambda_{v_{2}}(-k,k+p_{1}+p_{2},p_{3},p_{4})]D_{(l_{1})}(k^{2})D_{(l_{2})}((p_{1}+p_{2}+k)^{2})\end{split}$ (79)
Here, it is understood that the $\lambda_{v}(p,-p,-k,k)$ is fully renormalized and is decomposed to ones in channels $s,t,u$. The amplitude (79) is a weighted one-loop amplitude and is going to be a part of the effective quartic coupling. Recall that the subscripts $v$ without parenthesis indicate that the corresponding amplitude contains only the $v$-order contributions, respectively. Since each $\lambda_{v}$ and $m_{l}^{2}$ contains the $v$- and $l$-loop corrections, respectively, to obtain order-3 amplitudes, we need $v_{1}+v_{2}+l_{1}+l_{2}=3$. One problem is that we have no way to keep the order-3 part of the propagator and the propagator $D_{(3)}$ does not contain only the pure order-3 contribution, so we have a redundancy that we should remove.
We have five irregular diagrams in Fig. 7. They are all non-planar
$\begin{split}V_{3}^{\text{ir},1}({\underline{p}})&=(-6i\lambda)^{4}\int\frac{d^{4}k_{1}}{(2\pi)^{4}}\frac{d^{4}k_{2}}{(2\pi)^{4}}\frac{d^{4}k_{3}}{(2\pi)^{4}}D_{0}(k_{1}^{2})D_{0}((k_{1}+p_{1}+k_{2})^{2})D_{0}(k_{2}^{2})D_{0}(k_{3}^{2})\\ &\qquad\times D_{0}((k_{1}+p_{1}+p_{2}+k_{2}-k_{3})^{2})D_{0}((k_{1}+k_{3}-p_{4})^{2}),\end{split}$ (80)
$\begin{split}V_{3}^{\text{ir},2}({\underline{p}})&=(-6i\lambda)^{4}\int\frac{d^{4}k_{1}}{(2\pi)^{4}}\frac{d^{4}k_{2}}{(2\pi)^{4}}\frac{d^{4}k_{3}}{(2\pi)^{4}}D_{0}(k_{1}^{2})D_{0}(k_{2}^{2})D_{0}((p_{1}+p_{2}+k_{1}+k_{2})^{2})\\ &\qquad\times D_{0}(4p_{2}^{2})D_{0}((p_{1}+p_{3}+k_{2})^{2})D_{0}((p_{3}-p_{2}-k_{2})^{2}),\end{split}$ (81)
and three more permutations of the latter.
The total amplitude is the sum of all (79), (80) and (81). This can be compactly denoted as
$\begin{split}V_{3}(p_{1},p_{2},p_{3},p_{4})&=\frac{1}{2}\int\frac{d^{4}k}{(2\pi)^{4}}[-6i\lambda_{(2)}(p_{1},p_{2},k,-k-p_{1}-p_{2})]\\ &\qquad\times[-6i\lambda_{(2)}(-k,k+p_{1}+p_{2},p_{3},p_{4})]D_{(2)}(k^{2})D_{(2)}((p_{1}+p_{2}+k)^{2})\bigg{|}_{3}\\ &\quad+\sum_{i=1}^{5}V_{3}^{\text{irr},i}({\underline{p}}),\end{split}$ (82)
where $|_{3}$ extracts only the order-$3$ amplitude. Here in the last line, each coupling $\lambda_{(L-1)},m^{2}_{(L-1)}$ contains all the contributions from the zeroth order to ${\cal O}(\hbar^{L-2}\lambda^{L-1})$. Also, the order $L$ in the definition of $\overline{\Gamma}_{(L)}(p^{2})$ means that all the corrections up to $L$-loops are contained in the amplitude. In this, using arbitrary high loop-order couplings gives the same result; however, in practice, we should calculate the coupling to the possible lowest order. [^5] If there is no restriction $|_{L}$, we may have a more precise value for the amplitude than ${\cal O}(\hbar^{L+1}\lambda^{2L+2})$. In general, this amplitude gives a part of loop diagrams up to the $(2L)$-th loop. If we calculate order by order, we may omit calculation of the amplitude beyond the $(2L)$-th loop; however, if we already have a closed form on the quantities $\lambda_{(L)},m_{(L)}^{2}$, we obtain a slightly more precise value. It is pictorially displayed in Fig. 3. We expand all the vertices and lines with order-$L$ ones, and after the calculation, we extract the order-$L$ contributions. In practice, without this extraction, we have better accurate results with more accuracy.
Then we renormalize. Since all the component $\lambda_{v_{1}},\lambda_{v_{2}},D_{l_{1}},D_{l_{2}}$ are renormalized, we have no subdiagram divergence by construction. The irreducible amplitudes have no subdiagram divergences. The only divergence comes from the overall one-loop structure.
For the order-4 vertex, we make use of the previous result. The reducible parts (79) are obtained by replacing every vertex and line to $\lambda_{(3)}$ and $D_{(3)}$ and extracting the order-4 part. The formal structure of the amplitude is the same as that of the order-3 in (82), which is another aspect of self-similarity. Also, the order-3 irreducible amplitudes (80) and (81) are expanded to give order-4 amplitudes. This is depicted in Fig. 7.
This is to be applied to the two-point correlation function and the effective operator $\overline{\Gamma}^{(2)}$. The resulting propagator is also obtained as the inverse $\Delta({\underline{p}})=i/\overline{\Gamma}^{(2)}({\underline{p}})$ as in (58).
4.3 Overlapping divergence
The above generating method on higher-order couplings guarantees no subdiagram divergences. The prescription (50) also removes the overall divergence.
Figure 8: Overlapping divergence in a two-loop diagram in QED (top). It is automatically taken care of if we generate it (as the sum of two diagrams) using renormalized couplings $[-i.e.,_{(1)}({\underline{p}})]$ in the weighted one-loop vacuum polarization graph (bottom). The last two diagrams give the same amplitudes and are summed up to yield the correct multiplicity.
This should hold true if we happen to have, as a result of expansion, overlapping loops sharing an internal line. An overlapping diagram has more than one loop sharing an internal line, like the sunset diagram in Fig. 3, the third diagram on the second row in Fig. 10 and the one in Fig. 8. It has an ambiguous origin of divergence since the shared propagator carries the momenta of both loops.
In our iterative renormalization, the divergence from the overlapping loop is automatically taken care of in a symmetrical way. Consider a two-loop vacuum polarization diagram, the top panel in Fig. 8. We can obtain it by expanding a weight one vertex of the one-loop Feynman diagram as in the middle panel in Fig. 8, in which there are two cases. By construction, the expansion makes the subdiagram divergence-free, we only need to renormalize the overall part as in the lower panel in 8.
This is the consequence of the following theorem.
Theorem 7
(Bergère and Zuber [34]) In overlapping two loops, divergence comes from each loop separately, not the two or the whole at the same time.
We generate the two-loop vacuum polarization amplitudes by expanding the order-two photon propagator $\Pi^{\mu\nu}_{2}(p^{2})$. In turn, it is given by a weighted one-loop amplitude, in which all the internal propagators and vertices are of loop order one,
[…]
There are two places that one-loop vertex function $\gamma^{\mu}_{(1)}(p^{2})$, renormalized, finite and momentum-dependent, appear. They have further expansion giving the two diagrams in the middle panel in Fig. 8. Initially, we have not considered this; however, we correctly generated this diagram if we start from the order-two photon propagator $\Pi_{(2)}(p^{2})$.
Suppose we expand asymmetrically one of the vertices in the one-loop diagram, although it is easy to track the divergence. In that case, it is not clear whether the resulting amplitude becomes symmetric under the exchange of the two vertices. The two amplitudes obtained by expanding two $\gamma_{1}$’s are the same, since the corresponding diagrams are related by 180-degree rotation accompanied by $p\to-p$.
Extracting two-loop amplitude by expansion as in the middle panel in Fig. 8, we have
$\begin{split}-i\Pi^{\mu\nu}_{2*}(p^{2})&=-\frac{1}{2}\int\frac{d^{4}k}{(2\pi)^{4}}\operatorname{tr}[-ie\gamma_{(1)}^{\mu}(k,p+k)]S(k)[-ie\gamma_{(1)}^{\nu}(p+k,k)]S(p+k)\bigg{|}_{2}\\ &=\frac{e^{2}}{2}\int\frac{d^{4}k}{(2\pi)^{4}}\operatorname{tr}[\gamma^{\mu}+\gamma_{1}^{\mu}(k,p+k)]S(k)[\gamma^{\nu}+\gamma_{1}^{\mu}(k,p+k)]S(p+k)\bigg{|}_{2}\\ &=\frac{e^{2}}{2}\int\frac{d^{4}k}{(2\pi)^{4}}\operatorname{tr}\gamma^{\mu}\left[1+i\frac{e^{2}}{2}\int\frac{d^{4}k^{\prime}}{(2\pi)^{4}}\frac{1}{(k^{\prime}-k)^{2}}\big{(}S(k^{\prime}+p)-S(k^{\prime})\big{)}\gamma^{\rho}S(k^{\prime})\gamma_{\rho}\right]\\ &\quad\times S(k)\gamma^{\nu}\left[1+i\frac{e^{2}}{2}\int\frac{d^{4}k^{\prime}}{(2\pi)^{4}}\frac{1}{(k^{\prime}-k)^{2}}\big{(}S(k^{\prime}+p)-S(k^{\prime})\big{)}\gamma^{\rho}S(k^{\prime}+p)\gamma_{\rho}\right]S(k+p)\bigg{|}_{2},\end{split}$ (84)
where the trace acts on all the gamma matrices and the propagators on the right, using the standard notation [26]. Since ${\cal O}(e^{4})$ terms are the same, we add the two terms then the overall factor in front of the integral becomes $22\cdot e^{4}/4=e^{4}/$ 2. Since it, as a one-loop amplitude, has to be renormalized. Subtracting, we have
$\begin{split}i\overline{\Pi}^{\mu\nu}_{2}(p^{2})&=i\Pi^{\mu\nu}_{2}(p^{2})-i\Pi^{\mu\nu}_{2}(0)-ip^{2}\frac{d\Pi^{\mu\nu}_{2}}{d(p^{2})}(0)\\ &=\frac{ie^{4}}{2}\int\frac{d^{4}k}{(2\pi)^{4}}\frac{d^{4}k^{\prime}}{(2\pi)^{4}}\operatorname{tr}\gamma^{\mu}\frac{1}{(k^{\prime}-k)^{2}}\big{(}S(k^{\prime}+p)-S(k^{\prime})\big{)}\gamma^{\rho}S(k^{\prime})\gamma_{\rho}S(k)\gamma^{\nu}S(k+p)\\ &\ -\frac{ie^{4}}{2}p^{2}\int\frac{d^{4}k}{(2\pi)^{4}}\frac{d^{4}k^{\prime}}{(2\pi)^{4}}\operatorname{tr}\gamma^{\mu}\frac{1}{(k^{\prime}-k)^{2}}\frac{dS}{d(p^{2})}(k^{\prime})\gamma^{\rho}S(k^{\prime})\gamma_{\rho}S(k)\gamma^{\nu}S(k)\\ &\ -\frac{ie^{2}}{2}p^{2}\int\frac{d^{4}k}{(2\pi)^{4}}\operatorname{tr}\gamma^{\mu}S(k)\gamma^{\nu}\frac{dS}{d(p^{2})}(k).\end{split}$ (85)
Now, thanks to the correct factor,
$p^{2}\frac{dS(p^{2})}{d(p^{2})}=\frac{p_{\mu}\gamma^{\mu}}{2}\frac{dS(p^{2})}{d(p_{\mu}\gamma^{\mu})},$ (86)
we may regard (85) as the Taylor expansion of the following
$\begin{split}i\overline{\Pi}^{\mu\nu}_{2}(p^{2})&\simeq\frac{ie^{4}}{2}\int\frac{d^{4}k}{(2\pi)^{4}}\frac{d^{4}k^{\prime}}{(2\pi)^{4}}\operatorname{tr}\gamma^{\mu}\frac{1}{(k^{\prime}-k)^{2}}\big{(}S(k^{\prime}+p)-S(k^{\prime})\big{)}\gamma^{\rho}S(k^{\prime})\gamma_{\rho}S(k)\gamma^{\nu}S(k+p)\\ &\ -\frac{ie^{4}}{2}\int\frac{d^{4}k}{(2\pi)^{4}}\frac{d^{4}k^{\prime}}{(2\pi)^{4}}\operatorname{tr}\gamma^{\mu}\frac{1}{(k^{\prime}-k)^{2}}\big{(}S(k^{\prime}+p)-S(k^{\prime})\big{)}\gamma^{\rho}S(k^{\prime})\gamma_{\rho}S(k)\gamma^{\nu}S(k).\end{split}$ (87)
The first line is two-loop vacuum polarization, whose subdiagram divergences are canceled by the following counterterms. The next line removes the overall divergence. This makes clear how a renormalized coupling takes care of possible overlapping divergence. However, the form is symmetric under the exchange of the two vertices.
We also learn that the same diagram can be obtained by expansion from different diagrams. From the traditional viewpoint of renormalization, such different diagrams require different counterterms. Since the generation of a higher-loop diagram covers all the possible cases with renormalized subdiagrams, there is no subtlety of renormalization from the overlapping loop diagrams. In particular, the correct coefficients are automatically determined by combinatorics.
When we expand a vertex in the loop, we may generate an overlapping loop diagram. However, if the overlapping arises from the expansion of a vertex, the expansion gives the correct combinatorics. We can mechanically apply the expansion procedure.
If the overlapping arises from the expansion of a propagator, we correct the symmetric factor by multiplying that of the original diagram before expansion, followed by dividing the new diagram. This local amendment is sufficient to obtain the correct coefficients of the amplitude.
4.4 Finding irreducible amplitudes
An irreducible amplitude is an elementary unit of expansion. The order of an irreducible amplitude can be arbitrarily large, and we have no general method of generating it. In most cases, nevertheless, we can find ones according to the following observations.
Theorem 8
An irreducible diagram has
  1. 1.
    a subloop containing a line that connects non-adjacent vertices,
  2. 2.
    which makes overlapping loops.
We define two vertices are adjacent if there is no intermediate vertex on all the line connecting two vertices.
Consider a subamplitude of irreducible amplitude. It is non-renormalizable and has many external legs according to the Weinberg theorem. To make the total amplitude $r$-contractable, or equivalently, renormalizable, some external legs must be closed. This makes another loop that is overlapping, sharing this closed line. This new closed loop must not be contractable by the definition of the irreducibility of the total amplitude. An internal line connecting non-adjacent vertices makes one of the overlapping amplitudes non-contractable.
Note that the converse of the last statement is false: not every diagram with an internal line connecting non-adjacent vertices is irreducible. One counterexample is a two-loop overlapping diagram in $\phi^{3}$-theory, made with cubic interactions. If each loop contains two external lines, although the overlapping line connects non-adjacent vertices, each subdiagram is renormalizable, making the total amplitude reducible. This is a special case with the super-renormalizable operator. With the marginally renormalizable interactions, the internal line connecting non-adjacent vertices makes the resulting loop non-renormalizable.
From the above characterization of the line connecting non-adjacent vertices, we learn one helpful way to make irreducible diagrams. We can make it from a reducible loop diagram by rewiring the internal line.
Definition 6
By flop change, we exchange the ends of two different internal lines.
Examples are shown in Fig. 6, in which the left on each line is obtained from the right, which can be generated from tree-level coupling.
Another property of flop change is it preserves the number of loops.
Lemma 2
A flop change does not modify the numbers of vertices, lines and loops.
By definition, it is an action of changing the ending vertices of lines. Their numbers do not change. Thus, from the Euler characteristic relating the number of internal lines $I$ and the number of vertices $V$ [^6] A graph with $V$ vertices, $I$ edges and $L$ faces is regarded as a polygon projected on a surface thus the relation (88) is the Euler number of polyhedra with one face removed.
$L=I-V+1,$ (88)
the number of loops remains unchanged.
Most irreducible diagrams are generated by flop changes of other reducible diagrams of the same loop order. In fact, the Euler characteristic (88) is a very strict restriction for the existence of a loop diagram. Another restriction is the number of lines that cannot exceed the combination $(^{v}_{2})$ with the number of the vertices $v$. In a generic theory in the sense of Definition 3, almost all the irreducible diagrams are obtained. For a single $\phi^{4}$-theory or QED, the left diagrams are obtained by the flop changes of the right ones, respectively. All the right diagrams can be contracted to quartic coupling or photon-electron interaction. Therefore, it is a good search strategy first to generate an $L$-loop reducible diagram by expansions and perform the flop change.
However, not all the irreducible diagrams obtained by flop change. In a non-generic theory with a limited number of coupling, we cannot perform a flop change from an irreducible diagram to a reducible one. For instance, in the first line in Fig. 6, the irreducible diagram on the left cannot be obtained from the reducible diagram if the internal fields are different $a\neq b$. One good remedy is to make the Lagrangian generic by filling in missing operators, then make the diagram and perform a flop change. We can then check whether the irreducible diagrams obtained exist in the original theory.
5 Renormalization using weighted Feynman diagrams
So far, we have focused on generating amplitudes $\overline{\Gamma}_{L}^{(n)}$ that are fully renormalized. The problem we usually face is the reverse: how do we renormalize a given unrenormalized amplitude $A_{L}$? We only need to find the renormalized one $\overline{A}$ among the components $\overline{\Gamma}_{L}^{(n)}$. We develop a method of renormalization making use of it.
5.1 Renormalization as contraction-generation
We can renormalize an amplitude using contraction followed by generation. We contract all the subdiagrams, which corresponds to replacing loop-corrected parameters with uncorrected ones. Then, we perform the generation of the same order. As a result, we obtain a number of fully renormalized diagrams containing the original diagram.
Theorem 9
For an irreducible diagram, renormalization can be done by $r$-contraction and generation to the same order.
Remember that the generation gives us the completely renormalized coupling at the order. In this sense, contraction-generation is different from just removal of infinity because there are many terms in $\overline{\Gamma}^{(n)}_{L}$ other than $A^{(n)}_{L}$. However, to compare the calculation with the experiment, we need to compare $\overline{\Gamma}^{(n)}_{0}$ in $\overline{\Gamma}^{(n)}_{(L)}$. So, in this paper, we defined the renormalization in this complete sense.
On the other hand, $\overline{\Gamma}^{(n)}_{L}({\underline{p}},\underline{k_{i}})$ is a $c$-number function after evaluating the integral. We may also regard it as a (renormalized) coupling, with which we use the contracted Feynman diagram. In fact, this is useful in using an expanded diagram in which $A_{v}$ is a completely renormalized function. To distinguish it from the $r$-contracted diagram, we may assign the loop weight $v$ to this vertex.
Definition 7
A loop subamplitude of order $v$, $\overline{\Gamma}_{v},$ that is completely renormalized
$A_{v}\to\overline{\Gamma}_{v}.$ (89)
is called a vertex of weight $v$. It is called as replacement. We do the same to the lines.
We may also use the diagram (64), but we indicate the corresponding vertex using a big dot and denote the weight in the vicinity as in Fig. 7.
Definition 8
The replacement (89) gives a weighted Feynman diagram, in which some subdiagrams $\{\gamma_{v}\}$ corresponding to the replaced subamplitude is contracted $\Gamma/\{\gamma_{v}\}$. We indicate the order of $\overline{\Gamma}_{v}$ in the vicinity of the vertices. We do the same for the replaced lines.
Likewise, we keep the order $l$ of the propagator $\Delta_{l}$. Since we call a subamplitude a vertex and do not change the loop order, the total loop order (69) is invariant.
The weighted Feynman diagram has the same overall divergence structure as the unweighted Feynman diagram. Essentially, it is because there is no extra subdiagram divergence that occurs if we use effective couplings. This can be used as a part of the proof of Theorem 1 by Weinberg.
5.2 Method of renormalization
Given an amplitude of the loop order $L$, its renormalization is done by the reverse of the above generating process. Basically, we replace all the subdiagrams with the equivalent renormalized ones as in (89). By iteration, we arrive at a tree-level coupling of the loop order $L$ in the sense of (89), containing all the amplitude, including the original one.
Theorem 10
The following procedure is to generate, from an unrenormalized 1PI amplitude with the diagram $\Gamma$, a renormalized one.
  1. 1.
    We look for and contract an irreducible subdiagram $\gamma$. Suppose that this subdiagram is of “the $K$-loop order” in the sense of (69). Then, this subdiagram contracts to a point with $n$-point coupling
    $A^{(n)}_{K}(y,\gamma),$ (90)
    where $y$ is the location of the contracted $\gamma$ or the vertex in $\Gamma/\gamma$. For example, on each row in Fig. 4, one of the one-loop subdiagrams on the RHS is contracted to a point, yielding the diagram on the LHS.
  2. 2.
    For a propagator, we do the same. If we have a linear chain of identical 1PI diagrams, as in Fig. 2, we contract them at once to make one propagator as in (77),
    $A_{(K)}^{(2)}=D_{{(K)}}^{-1}(y,\gamma),$ (91)
    which is also recorded as one entry. In this case, the order $K$ is the order of one 1PI component.
  3. 3.
    We obtain the weighed Feynman diagram. Record the resulting coupling in the “stack”, that is, a sequence forming an ordered set
    ${\cal P}=\left(A^{(n_{1})}_{K_{1}}(y_{1},\gamma_{1}),A^{(n_{2})}_{K_{2}}(y_{2},\gamma_{2}),A^{(n_{3})}_{K_{3}}(y_{3},\gamma_{3}),\dots,A^{(n_{N})}_{K_{N}}(y_{N},\gamma_{N})\right),$ (92)
    where the subscript means the order.
  4. 4.
    In the end, we obtain a tree diagram of the total loop order
    $\sum_{i=1}^{N}K_{i}.$ (93)
  5. 5.
    Then we obtain a series of nested renormalized diagrams backward: we retrieve the record from the previous set $\cal P$ in the reverse order and generate the renormalized loop subdiagrams $\{\overline{\Gamma}_{K_{i}}(y_{i},\gamma_{i})\}$ at the desired place $y_{i}$ by the expansion,
    $\gamma({\cal P}^{-1})=\big{(}\overline{\Gamma}_{K_{N}}(y_{N},\gamma_{N}),\dots,\overline{\Gamma}_{K_{3}}(y_{3},\gamma_{3}),\overline{\Gamma}_{K_{2}}(y_{2},\gamma_{2}),\dots,\overline{\Gamma}_{K_{1}}(y_{1},\gamma_{1})\big{)}.$ (94)
    This takes us back to the original diagram now with the overall divergence and renormalized parameters $\overline{\Gamma}$.
There can be more than one way to perform the contraction. They commute and we arrive at the same weighted Feynman diagram, containing no loop. If we generate the renormalized amplitude containing the original one, following the procedure 5, we get the complete effective coupling with the correct coefficients. However, in doing Procedure 5, the expansion, if we generate along all the paths yielding the original diagram, amounts to taking the finite part of the original diagram.
The expansion of effective couplings of a higher order to loop diagrams containing lower order couplings may give rise to a number of the same diagrams via different paths. For instance, both diagrams on the right in Fig. 9 arrive at the diagram in the middle. Indeed, the original diagram $\Gamma$ can be contracted in a number of ways, giving different weighted diagrams. Conversely, these different diagrams expand to the same diagram $\Gamma$. The generating method tells us that we have to consider all of these possibilities to have the correct multiplicity, which is the same as that of the original diagram $\Gamma$ from the symmetry consideration.
5.3 An example
Figure 9: Renormalization of a 15-loop diagram. Taking into account 1-particle-reducible (1PR) subdiagrams in the propagators, the effective loop order is 13. After contracting all the subdiagrams, indeed, we obtain an order-13 propagator. Each effective couplings are renormalized.
We take an example in the $\phi^{4}$-theory. Consider an amplitude with the 15-loop Feynman diagram given in Fig. 9. Taking into account one-particle-reducible subgraphs in the propagators, it is effectively a 13-loop diagram; since it has two scalar external legs, it reduces to a scalar propagator. So, we expect that this appears as one of the graphs in the propagator $D_{(13)}(p^{2})$.
For this particular diagram, the renormalization procedure is as follows. First, we have 1PR parts in (a) and (b), each of which belongs to an order one propagator $D_{(1)}({\rm a})$ and $D_{(1)}({\rm b}),$ respectively. The subdiagram (c) is an irreducible four-loop order with four external legs and belongs to an order four quartic coupling $\lambda_{4}({\rm c})$.
Contracting all of these, we get the second diagram in Fig. 9. The part (d) subdiagram, a one-loop-corrected propagator containing $D_{(1)}({\rm a})$, becomes the part (f) subdiagram in the last diagram $D_{(2)}({\rm d})$, after contracting the loop. The subdiagram (e) is again an irreducible diagram of order 4, containing the order four coupling $\lambda_{4}({\rm c}^{\prime})$. Therefore, with the order four vertex in the 4-loop diagram, we have an order $4+4=8$ vertex $\lambda_{8}({\rm e})$ as in the last diagram.
Contracting all, we obtain the last diagram that has the same structure as the sunset diagram. We have an order two propagator and an order one propagator in part order 11 vertex, forming a 2-loop diagram. Contracting fish diagrams twice, we obtain an order-13 propagator $D_{(13)}$.
Alternatively, from the middle diagram, we may take another contraction to obtain a different effective coupling. For example, we may contact the subdiagram (d) first. The upper loop makes an effective propagator of order 2. The resulting quartic coupling has the order $1+2+4+4=11$. With this, the final diagram is another sunset (2-loop) diagram with unmodified propagators. It can be contracted to make the order 13-diagram for $D_{(13)}(p^{2})$. The above process is many-to-one, arriving at the same final diagram.
The first records the paths
$\left(D_{(1)}({\rm a}),D_{(1)}({\rm b}),\lambda_{4}({\rm c}),D_{(2)}({\rm d}),\lambda_{4}(\rm{c}^{\prime}),D_{(12)},D_{(13)}\right).$
Reading from the last, we expand the propagators and the couplings. We expand $D_{(13)}$ by a series of order one expansion $D_{(13)},D_{(12)}$. Then, we expand the order eight vertex on the right by expanding the order four irreducible diagram. The rest is straightforward, and we obtain the completely renormalized amplitude.
5.4 The forest formula
We have done the bottom-up renormalization for a general amplitude. However, it is useful to compare it with the traditional one, the top-down method, using the recursion relation (55). Zimmermann found the set of “forests” below solves the recursion relation (55) [17].
Theorem 11
(Zimmermann [17]) For the amplitude $A$ with the graph $\Gamma$, we obtain the renormalization $\overline{A}$ by
$\overline{A}=(1-t^{\Gamma})\sum_{\{{\cal F}\}}\bigg{[}\prod_{\gamma\in{\cal F}}(-t^{\gamma})\bigg{]}A,$ (95)
where the forest ${\cal F}$ is a set of disjoint subgraphs of $\Gamma$, including the empty set, some of which can be nested in another.
Here, the formula (95) is equai to the original relation (55): Using Theorem 7, we can remove the contributions from the disjoint subamplitude $(1-t^{\gamma_{i}})(1-t^{\gamma_{j}})=1$ if $\gamma_{i}\cap\gamma_{j}\neq\emptyset$.
Since our generated amplitude $\overline{\Gamma}^{(n)}_{L}$ is fully renormalized, we may extract the forest solving the relation (95).
Definition 9
Let the renormalization of an amplitude $A^{(n)}_{L}$ be $\overline{A}^{(n)}_{L}$. This $\overline{A}^{(n)}_{L}$ contains the Taylor expansion of the subamplitudes $\gamma$, that is, a number of terms acted by the product of Taylor operators $\prod t^{\gamma}$. For each term, we contract the corresponding subgraph $\gamma$ when applying $t^{\gamma}$. We define the forest ${\cal F}$ as the set containing the set of graphs undergoing the contractions simultaneously by $\prod t^{\gamma}$.
Since we are only interested in the divergence structure, we abbreviate a general amplitude (59) with the corresponding Feynman graph $\gamma$ as
$A=\int_{\gamma}\Delta_{\gamma}g_{\gamma},$ (96)
with the total product of the propagators $\Delta_{\gamma}$ from all the lines in $\gamma$ and the total product of the couplings $g_{\gamma}$ from all the vertices in $\gamma$ [^7] The integral can be omitted in our discussion. We suppressed the momentum dependence.
First, we have nesting diagrams, defining the subgraph. Consider a diagram $\gamma$ and its subdiagram $\gamma_{i}\subset\gamma$. The amplitude is
$A=\int_{\gamma/\gamma_{i}}\Delta_{\gamma/\gamma_{i}}g_{\gamma/\gamma_{i}}\int_{\gamma_{i}}\Delta_{\gamma_{i}}g_{\gamma_{i}}$ (97)
Further contractions are possible for further nesting structures. Performing subamplitude renormalization on $\gamma_{i}$, we have
$\int_{\gamma/\gamma_{i}}\Delta_{\gamma/\gamma_{i}}g_{\gamma/\gamma_{i}}\int_{\gamma_{i}}[1-t^{\gamma_{i}}]\Delta_{\gamma_{i}}g_{\gamma_{i}}$ (98)
The full renormalization of the amplitude is
$\begin{split}\overline{A}&=[1-t^{\gamma}]\int_{\gamma/\gamma_{i}}\Delta_{\gamma/\gamma_{i}}g_{\gamma/\gamma_{i}}\int_{\gamma_{i}}[1-t^{\gamma_{i}}]\Delta_{\gamma_{i}}g_{\gamma_{i}}\\ &=\int_{\gamma/\gamma_{i}}\Delta_{\gamma/\gamma_{i}}g_{\gamma/\gamma_{i}}\int_{\gamma_{i}}\Delta_{\gamma_{i}}g_{\gamma_{i}}-t^{\gamma_{i}}\int_{\gamma/\gamma_{i}}\Delta_{\gamma/\gamma_{i}}g_{\gamma/\gamma_{i}}\int_{\gamma_{i}}\Delta_{\gamma_{i}}g_{\gamma_{i}}\\ &-t^{\gamma}\int_{\gamma/\gamma_{i}}\Delta_{\gamma/\gamma_{i}}g_{\gamma/\gamma_{i}}\int_{\gamma_{i}}\Delta_{\gamma_{i}}g_{\gamma_{i}}+t^{\gamma}t^{\gamma_{i}}\int_{\gamma/\gamma_{i}}\Delta_{\gamma/\gamma_{i}}g_{\gamma/\gamma_{i}}\int_{\gamma_{i}}\Delta_{\gamma_{i}}g_{\gamma_{i}}.\end{split}$ (99)
The first term in the second line is the original $A$. Note that the $t$-operator on the right acts first. Now, contract the part of the graph if it is acted by the Taylor expansion operator $t$. The $t$-acted part is the counterterm and we may indicate it in the diagram. Reading from the amplitude gives the forests [^8] Thus, we have four forests. We omit their labels.
${\cal F}_{\text{nes}}=\{\gamma,\gamma_{i}\},\{\gamma\},\{\gamma_{i}\},\emptyset.$ (100)
The inductive method of generating higher loop diagrams using effective couplings corresponds to considering nested forests of subdiagrams. If a coupling belongs to a higher-order coupling, the corresponding lower-order subdiagram is a part of the higher-order (sub)diagram. We say that the former is nested in the latter.
Next, we consider the product amplitude from disjoint diagrams $\gamma_{i},\gamma_{j},\gamma_{i}\cap\gamma_{j},$. We have
$A=\int_{\gamma_{i}}\Delta_{\gamma_{i}}g_{\gamma_{i}}\int_{\gamma_{j}}\Delta_{\gamma_{j}}g_{\gamma_{j}}.$ (101)
They make independent coupling when contracted so that they can be separately renormalized
$\begin{split}\overline{A}&=[1-t^{\gamma_{i}}]\int_{\gamma_{i}}\Delta_{\gamma_{i}}g_{\gamma_{i}}[1-t^{\gamma_{j}}]\int_{\gamma_{j}}\Delta_{\gamma_{j}}g_{\gamma_{j}}\\ &=\int_{\gamma_{i}}\Delta_{\gamma_{i}}g_{\gamma_{i}}\int_{\gamma_{j}}\Delta_{\gamma_{j}}g_{\gamma_{j}}-t^{\gamma_{i}}\int_{\gamma_{i}}\Delta_{\gamma_{i}}g_{\gamma_{i}}\int_{\gamma_{j}}\Delta_{\gamma_{j}}g_{\gamma_{j}}\\ &-\int_{\gamma_{i}}\Delta_{\gamma_{i}}g_{\gamma_{i}}t^{\gamma_{j}}\int_{\gamma_{j}}\Delta_{\gamma_{j}}g_{\gamma_{j}}+t^{\gamma_{i}}\int_{\gamma_{i}}\Delta_{\gamma_{i}}g_{\gamma_{i}}t^{\gamma_{j}}\int_{\gamma_{j}}\Delta_{\gamma_{j}}g_{\gamma_{j}}\\ \end{split}$ (102)
As before, after discarding, we have the forest
${\cal F}_{\text{disj}}=\{\gamma_{i},\gamma_{j}\},\{\gamma_{i}\},\{\gamma_{j}\},\emptyset.$ (103)
The collection of the forests has the structure of a power set,
$\{{\cal F}_{\text{nes}}\}=2^{\{\gamma,\gamma_{i}\}},\quad\{{\cal F}_{\text{disj}}\}=2^{\{\gamma_{i},\gamma_{j}\}},$ (104)
which is obvious from the subtraction structure in the effective coupling and its expansion. The relation holds if the diagrams are subamplitudes of a bigger diagram.
We know that an amplitude $A^{(n)}_{L}$ contains either nested or disjoint subdiagrams from the generation. Therefore, reversing the logic, we can prove the above relation from the forest. Our method of renormalization gives an equivalent result to the BPHZ scheme [14, 15, 16, 17]. The idea is the same: replace every divergent (sub)diagram with a renormalized one [14], which is justified by the addition of counterterms or the removal of the bare parameter. From our approach, the reason for using the forest is clear (See also [35, 36]).
We consider two more nontrivial cases. First, consider two disjoint graphs $\gamma_{i},\gamma_{j}$ contained. Applying the above procedures, we get the forests
$\{{\cal F}_{\text{all inside}}\}=2^{\{\gamma,\gamma_{i},\gamma_{j}\}},\quad\gamma_{i}\cap\gamma_{j}=\emptyset,\quad\gamma_{i}\cup\gamma_{j}\subset\gamma.$ (105)
Since $\gamma_{j}\subset\gamma$, the Taylor operators do not commute
$t^{\gamma}t^{\gamma_{j}}\neq t^{\gamma_{j}}t^{\gamma},$ (106)
ant the LHS is in the correct order. It should be understood that for $\{\gamma,\gamma_{j}\}$, we should apply the Taylor operator as $t^{\gamma}t^{\gamma_{i}}$ from the nesting structure. This is automatically done if we regard every subset in $\cal F$ by the nested ordering. The next case is when only $\gamma_{i}$ is nested in $\gamma$ and $\gamma_{j}$ is disjoint to both, we have
$\{{\cal F}_{\gamma^{j}{outside}}\}=2^{\{\gamma,\gamma_{i},\gamma_{j}\}},\quad\gamma\cap\gamma_{j}=\emptyset,\quad\gamma_{i}\subset\gamma.$ (107)
Formally, the answer looks the same. This time, the commutation relation holds
$t^{\gamma}t^{\gamma_{j}}=t^{\gamma_{j}}t^{\gamma}.$ (108)
Nevertheless, if we obey the ordering of the nested structure, we can apply the Taylor expansion without subtlety.
Finally, a given amplitude $A^{(n)}_{L}$ may have overlapping loops in general so that we have more than one way of defining subamplitudes. As we have seen in the generating method, taking all of the possible “routes” should give the correct coefficient for the renormalized one $\overline{A}^{(n)}_{L}.$ Therefore, we include all the routes in the forest.
This reasoning led us to the following theorem.
Theorem 12
The $\Gamma$-forest ${\cal F}$ is the nesting-ordered power set of disjoint (sub)diagrams $\gamma_{i}\subseteq\Gamma$ with $\delta(\gamma_{i})\geq 0$
{}$\{{\cal F}\}=\bigcup_{\text{routes}}{\cal N}\big{(}2^{\bigcup\{\gamma_{i}\}}\big{)},$ (109)
where ${\cal N}$ is the nesting-ordering operator.
Although the solution is equivalent, there is a little difference. As we have seen in Section 2, we can track the origin of the overlapping divergence. From the example of the sunset diagram, we find that we do not care about the overlapping and just need to modify the symmetric factor. This is consistent with Theorem 7, and we do not need to worry about overlapping in applying the forest formula.
Also, technically, our method is easier to use. In the forest formula, we need to find maximal subamplitudes. It is easier to find a small loop than to find a maximal subloop required in the forest formula. After contraction, the corresponding weighted Feynman diagram becomes simpler, on which we iterate the procedures.
6 Discussion
We presented a new (and old) view of renormalization that makes use of the self-similarity of amplitudes. A loop amplitude at any order $\overline{\Gamma}_{L}^{(n)}({\underline{p}})$ that is fully renormalized has self-similarity in the sense that it also plays the role of a coupling inside another amplitude or even itself. This means the sum of these in various loop order acts as an effective coupling, similar to the coupling in a free field theory.
Such coupling is best described by the effective couplings $\overline{\Gamma}^{(n)}({\underline{p}})$, the coefficients of the effective potential, including perturbative corrections. They are “physical” because they appear in the $S$-matrix and with them, we can test the theory in the experiments. The success of QFT is its prediction on the scale dependence of $\overline{\Gamma}^{(n)}({\underline{p}})$, in which the scale is specified by the set of the external momenta ${\underline{p}}$. In other words, we obtain how the correction changes the value of effectively the same coupling across scales. Loop correction modifies the coupling whose relevance depends on the scale ${\underline{p}}$. In renormalizable theory, a renormalized quantity has the subtraction (Taylor expansion) structure so that it is finite, independent of regularization and insensitive to cutoff, provided decoupling takes place. Otherwise, the unrenormalized amplitude is divergent and cannot become a small coupling admitting perturbation.
Having a well-defined unit of physical parameters as the fully renormalized effective couplings, the following features of QFT are clearly seen in the effective field theory. First, all the renormalization we use is finite, relating two couplings across a finite range as in (52) and the cancellation of divergence is a bi-product. The issue of infinity only lies in the bare parameter in the Lagrangian. Every quantity appearing in the effective action can be written in terms of observable instead of the bare ones, as Theorem 6 shows. Also, we use relative physical parameter $\overline{\Gamma}^{(n)}_{(L)}$(p) from the reference parameter difined at a scale. The scale is specified by the external momenta ${\underline{p}}$ and it should be kept in the effective action to renormalize the couplings. This is the prediction of QFT.
With this, we can systematically generate higher-loop corrected amplitudes that are fully renormalized. We can also prove the renormalizability of a given diagram, which is equivalent but an alternative description to BPHZ.
After renewing the concept of renormalization using the external momentum, we performed complete renormalization. It is not merely extracting the finite part but obtaining physical quantity. To do this, we need the sum $\overline{\Gamma}_{(L)}^{(n)}({\underline{p}})$ of complete couplings belonging to the same $n$-point correlation function up to the order $L$ that we want to calculate. It is because the renormalization (75) condition to be fixed by to the observed quantity at the reference scale, $\overline{\Gamma}_{0}^{(n)}({\underline{p}}),$ needs the full summation of couplings to the order we can access.
The generation method helps us understand overlapping divergences. An amplitude having overlapping divergence can be made asymmetrical by contracting to a loop amplitude without overlapping divergence. This explains the result of Bergère and Zuber, Theorem 7, and the reason why, in the BPHZ renormalization, we should make a forest without overlapping diagrams. A drawback is we have to recover the symmetric factor if the diagram after expansion acquires enhanced symmetry.
The uniqueness of the effective coupling in a given theory tells us what physical quantity we should compare with the experiment. Saving, retrieving, and sharing the results are possible, and we can do fully/partly numerical calculations.
Although our framework provides a good alternative way to solve the Bogoliubov recursion relation for renormalization, rather than performing the renormalization procedure, it is more appropriate to generate the diagram, giving the process that we are interested in. In other words, any given $L$-loop diagram is obtained from the order-$L$ coupling with the same external legs. In practice, we are not interested in calculating particular amplitude, but we want the sum of all diagrams of a given order.
However, we still lack a general method to generate all the irreducible diagrams exhaustively. The elements of expansion and generations are irreducible amplitudes. They are present at sufficiently high orders, but their patterns seem irregular. For this, we have provided a search method using flop change, however, it is not efficient. Understanding algebra behind Feynman diagrams would help [35, 36].
Figure 10: Up to two-loop diagrams are contained in the two-loop corrected propagator. In particular, if we replace the parameters of one-loop parameters, we obtain all the two-loop diagrams. In general, all the diagrams up to $n$-loop are contained in the $n$-loop propagator and in the one-loop diagram made of $(n-1)$-loop-corrected parameters.
Acknowledgments
This work is partly supported by the grant RS-2023-00277184 of the National Research Foundation of Korea. The author is grateful to Hyun-Min Lee for the discussions, to Michael Ratz, Yuri Shirman and Timothy Tait for the comments and to the Particle Physics Group of UC Irvine and the Theory Group of Chung-Ang University and hospitality while the work was finalized.